Non-Standard Operations

This list is intended to cover all non-standard operations of EFW, planned or accidental, from 2001-02-01.

Start dt (sec) Cluster ID Description Resources Plan short Desc
2001-02-23 05:18:00 11700 1 Very bad data due to high density. no info
2001-02-23 05:27:00 11160 2 Very bad data due to high density. no info
2001-02-25 23:26:00 2640 1 Bad timing of eclipse bias current change. no bad_data
2001-02-26 06:10:00 12960 1 2 Very bad data due to high density. no info
2001-03-05 02:42:00 2520 1 Eclipse. yes bad_data
2001-03-09 20:58:00 2340 1 Eclipse. yes bad_data
2001-03-12 12:16:00 9240 1 2 Very bad data due to high density. no info
2001-03-16 11:59:00 14160 1 Very bad data due to high density. no info
2001-03-16 11:59:00 18540 2 Very bad data due to high density. no info
2001-03-17 09:20:00 13260 1 2 Very bad data due to high density. no info
2001-03-19 00:18:00 10920 1 2 Very bad data due to high density. no info
2001-03-22 02:48:00 3120 1 2 Very bad data due to high density. no info
2001-03-22 03:40:00 99660 2 Miscommanded burst caused NULL sampling mode.
1. Anders 010322
2. Anders 010323
no no_tm
2001-03-22 03:40:00 10260 1 Miscommanded burst caused NULL sampling mode.
1. Anders 010322
2. Anders 010323
no no_tm
2001-03-22 03:40:00 113460 4 Miscommanded burst caused NULL sampling mode.
1. Anders 010322
2. Anders 010323
no no_tm
2001-03-22 03:40:00 118740 3 Miscommanded burst caused NULL sampling mode.
1. Anders 010322
2. Anders 010323
no no_tm
2001-03-23 20:06:00 6240 2 Very bad data due to high density. no info
2001-03-31 06:35:00 1380 1 2 3 4 Very bad data due to high density. no info
2001-04-12 11:00:00 60600 4 HXONLY sampling mode. no hxonly
2001-04-18 07:30:00 300 1 Atitude manoeuvre. yes bad_data
2001-04-18 09:00:00 300 2 Atitude manoeuvre. yes bad_data
2001-04-18 12:00:00 240 3 Atitude manoeuvre. yes bad_data
2001-04-21 21:30:00 11100 1 3 4 Very bad data due to high density. no info
2001-05-10 18:04:20 70 1 Axial delta-v manoeuvre. yes bad_data
2001-05-13 11:13:00 4020 3 Radial delta-v manoeuvre. yes bad_data
2001-05-13 14:28:00 2220 2 Radial delta-v manoeuvre. yes bad_data
2001-05-16 22:46:00 1680 1 Radial delta-v manoeuvre. yes bad_data
2001-05-17 12:56:00 360 2 Radial delta-v manoeuvre. yes bad_data
2001-05-17 17:27:00 2100 3 Radial delta-v manoeuvre. yes bad_data
2001-05-21 23:51:00 180 2 Axial delta-v manoeuvre. yes bad_data
2001-05-25 02:11:30 80 3 Axial delta-v manoeuvre. yes bad_data
2001-07-15 22:59:00 357360 4 Numerous problems as explained in resource e-mails.
1. Per-Arne 010727
2. Keith 010727
no bad_data
2001-07-16 21:45:00 240 3 Attitude manoeuvre. yes bad_data
2001-07-17 01:00:00 240 2 Attitude manoeuvre. yes bad_data
2001-07-17 07:15:00 260 1 Attitude manoeuvre. yes bad_data
2001-07-23 13:54:17 -1 2 The 10Hz filter on probe 3 failed at this time, and has never recovered. no no_10Hz_filt
2001-07-30 02:55:00 3600 4 HXONLY sampling mode. no hxonly
2001-08-15 03:07:03 2 2 Single bad HX packet. no bad_tm
2001-09-05 13:52:00 4320 2 Eclipse. yes bad_data
2001-10-01 07:30:00 2100 1 2 3 4 Verify sweep interval settings. Will cause loss of spin fits for 25 minutes. 07:30 After reset at start of AI, do not uplink spin patch or set burst. Start sweeps every 30 s (SC1), 1 min (SC2), 90 s (SC3), 2 min (SC4). 07:45 Setup burst (SWECJ349). 07:55 Spinfit patch (SWECJ354). 08:05 Test completed: change sweep interval to 30 mins on all. (SWECJ340 and 339).
1. Ieuan
yes no_spin_fits
2001-10-11 12:34:41 1258 1 TM loss due to s/c anomaly
1. ESOC 011012
2. ESOC 011015
no no_tm
2001-10-12 22:39:41 2 3 Single bad HX packet. no bad_tm
2001-10-15 01:18:00 61030 1 Probe 1 latched at -68 V, other probes also saturated much of the time, reason unknown.
1. Anders 011015
no bad_data
2001-10-17 05:48:43 90 1 2 3 4 Time tagged burst operations with 2E+3B components at perigee. Loss of data for a little more than a minute expected.
1. Commented excerpt from PIOR
2. Mail to Nicole and Milan
yes no_tm
2001-10-22 19:30:00 300 4 Attitude manoeuvre. yes bad_data
2001-10-22 21:15:00 300 3 Attitude manoeuvre. yes bad_data
2001-11-06 00:47:00 4980 2 Bad telemetry, reason unknown. no bad_tm
2001-11-07 19:59:00 2640 3 SC3 system reboot. No data during this time.
1. ESOC 011107
2. ESOC 011108
yes no_tm
2001-11-13 00:20:00 65400 2 Strange data for unknown reason. EDI high current? no bad_hx
2001-11-14 18:16:00 26880 1 Data loss due to ESA ground station problem
1. ESOC 011121
no no_tm
2001-11-18 06:07:14 39496 2 Probe 1 latched at +68 V. Probe 1 problem cured by EFW reset.
1. Per-Arne 011120
2. Anders 011129
no no_p1
2001-12-11 14:27:00 7020 1 2 3 4 Bad WEC telemetry, DWP buffer overflow? no bad_tm
2001-12-13 23:14:00 7020 1 2 3 4 Bad WEC telemetry, DWP buffer overflow? no bad_tm
2001-12-28 03:02:57 -1 1 Probe 1 effectively lost from this time.
1. Per-Arne 020102
2. Per-Arne 020104
3. Mats 020108
no no_p1
2001-12-29 20:33:00 7140 4 Bad WEC telemetry, DWP buffer overflow? no bad_tm
2002-01-02 06:58:00 17040 2 S/c problem caused data loss
1. ESOC 020108
no no_tm
2002-01-02 13:43:00 9480 2 S/c problem caused data loss
1. ESOC 020108
no no_tm
2002-01-02 14:25:00 4500 1 Attempts to bring P1 back to working order
1. Per-Arne 020102
no spec_bias
2002-01-06 08:26:40 240 1 Maneuvre? no bad_data
2002-01-06 17:04:00 650 4 Maneuvre? no bad_data
2002-01-07 17:36:00 77370 3 WEC latchup. All EFW data lost until 2002-01-08 15:05:30. WEC restarted, EFW data regained
1. Keith 020108
no no_tm
2002-01-08 09:50:00 -157 1 P2 bias went to 0xB2 (-300 nA), causing saturation at -68 V; later to 0 bias. Probable explanation is timing conflict between regular sweep command and extra bias command for probe 1.
1. Per-Arne 010209
2. Per-Arne 010209
no no_p2
2002-01-11 11:45:00 230 2 Attitude manoeuvre
1. Per-Arne 060330
yes bad_data
2002-01-11 14:30:00 180 3 Attitude manoeuvre
1. Per-Arne 060330
yes bad_data
2002-01-11 17:30:21 8 4 Spin-up manoeuvre
1. Per-Arne 060330
yes bad_data
2002-01-11 19:59:53 11 1 Spin-up manoeuvre
1. Per-Arne 060330
yes bad_data
2002-01-18 01:31:00 39660 1 WEC data lost following DWP watchdog reset on SC1.
1. Per-Arne 020118
2. Keith 020118
no no_tm
2002-02-13 03:37:48 3 1 Maneuvre? no bad_data
2002-02-15 18:55:00 604800 1 P2 in bias voltage (Langmuir) mode for a week from this time
1. Anders 020125
yes no_p2
2002-02-17 21:45:00 1500 1 2 3 4 Bias current +16 nA during perigee passes.
1. Anders 020125
yes spec_bias
2002-02-17 22:00:00 120 3 Test different bias currents at perigee eclipse.
1. Anders 020125
2. Anders 020125
yes spec_bias
2002-02-20 06:43:00 2100 1 2 3 4 Bias current +16 nA during perigee passes.
1. Anders 020125
yes spec_bias
2002-02-22 15:43:00 -157 1 2 3 4 Bias current +16 nA during perigee passes.
1. Anders 020125
yes spec_bias
2002-03-20 09:59:00 85440 2 CTU switchover on SC2 caused loss of all science data.
1. ESOC 020320
2. ESOC 020321
yes no_tm
2002-04-16 09:00:00 240 3 Attitude manoeuvre
1. Per-Arne 060331
yes bad_data
2002-04-16 13:00:00 240 4 Attitude manoeuvre
1. Per-Arne 060331
yes bad_data
2002-04-29 09:55:00 772560 3 WEC science telemetry bit error capable of causing all sorts of problems: timing jumps, apparent switches in EFW burst data, glitches.
1. Keith 020612
no bad_tm
2002-05-08 08:31:00 2280 3 WEC power cycled to clear up DWP memory latchup
1. Keith 020507
2. Silvano 020508
yes no_tm
2002-05-12 03:53:57 1800 1 2 3 4 Bias test, running -70 nA to all spheres
1. Anders 020423
2. Anders 020423
yes spec_bias
2002-05-26 10:27:00 54180 3 Corrupted WEC TM triggered s/c to reset WEC 020526 10:27. EFW then operated with default settings (e.g. non-ideal bias) until 020527 01:30.
1. Per-Arne 020527
2. Silvano 020527
yes spec_bias
2002-06-14 15:02:04 2 4 Single bad EFW TM packet
1. Per-Arne 060322
no bad_tm
2002-06-19 09:42:00 3120 3 Spacecraft manoeuvre contaminates data. no bad_data
2002-06-29 05:05:00 180 4 EFW telemetry completely confused.
1. Silvano 020704
2. Anders 020704
3. Keith 020710
no no_tm
2002-07-29 09:06:59 -1 3 Probe 1 effectively lost from this time.
1. Per-Arne 020806
no no_p1
2002-08-06 14:10:00 1200 2 3 WBD noise test
1. Per-Arne 060323
yes bad_tm
2002-08-06 18:20:00 10800 1 2 3 4 BM1 data intermittent (period about 20 s ???) due to wrong STAFF SA mode (TM overflow).
1. Keith 020808
2. Per-Arne 020821
no bad_tm
2002-08-10 05:34:13 194147 2 Latchup, SEU protection on ADC trigged, causing complete loss of data until power on/off cycle.
1. Per-Arne 020812
no bad_tm
2002-08-13 01:47:00 49560 2 Strange bias sweep behaviour: sweeps take longer and longer time, the last one (at 15:19:20) takes 45 s. They are flagged as sweeps or playbacks only for the nominal time (1-2 s), and hence some HX data are contaminated
1. Per-Arne 021217
no bad_tm
2002-08-16 01:10:00 51300 3 SC3 science and HK data lost due to s/c anomaly
1. Silvano 020823
no no_tm
2002-08-29 22:52:00 59880 2 SC2 reebooted by mistake at 22:5WEC not fully recovered until the end of a BM1 period at 2002-08-30 15:30.
1. ESOC 020830
2. Silvano 020830
3. Keith 020902
no no_tm
2002-08-30 03:30:00 8400 4 Data loss for s/c reasons.
1. ESOC 020914
no no_tm
2002-08-30 14:56:00 19380 4 Data loss for s/c reasons.
1. ESOC 020914
no no_tm
2002-08-30 21:08:00 1920 4 Data loss for s/c reasons.
1. ESOC 020914
no no_tm
2002-08-30 23:00:00 3600 4 Data loss for s/c reasons.
1. ESOC 020914
no no_tm
2002-08-31 00:00:00 7200 4 Data loss for s/c reasons.
1. ESOC 020914
no no_tm
2002-08-31 00:02:12 7680 4 Data loss for s/c reasons.
1. ESOC 020914
no no_tm
2002-09-21 01:18:00 11340 1 S/c problem caused loss of sun reference pulse and thus scrambling of WEC data. The EFW onboard spinfits should be crazy, as there is no sun pulse.
1. Silvano 020922
2. Keith 020923
no bad_tm
2002-10-02 00:49:00 37080 4 EFW data scrambled, though analog signals to WHISPER and STAFF are all right. Presumably a radiation induced SEU.
1. Per-Arne 021002
2. Silvano 021004
3. Per-Arne 021006
no no_tm
2002-10-02 16:00:00 34200 1 2 3 4 Testing the FGM burst trig flag
1. Anders 020910
2. Anders 020910
3. Anders 021104
yes info
2002-10-05 02:30:00 23400 1 2 3 4 Testing the FGM burst trig flag
1. Anders 020910
2. Anders 020910
3. Anders 021104
yes info
2002-10-07 19:00:00 300 2 Attitude manoeuvre yes bad_data
2002-10-07 20:00:00 300 4 Attitude manoeuvre yes bad_data
2002-10-07 22:45:00 300 1 Attitude manoeuvre yes bad_data
2002-10-10 14:07:00 77460 3 WEC turned off by SC after DWP voltage flag = 0xFF.
1. Silvano 021011
2. Silvano 021011
no no_tm
2002-10-20 00:52:28 746 3 Ground station problems caused data loss.
1. ESOC 021022
no no_tm
2002-10-20 13:21:52 75345 4 WEC turned off by s/c
1. Silvano 021021
2. Keith 021021
3. Silvano 021021
4. Keith 021021
no no_tm
2002-10-21 02:32:17 4106 3 Ground station problems caused data loss.
1. ESOC 021022
no no_tm
2002-10-28 03:00:00 -157 2 4 Two-minute bias current sweeps using internal burst at 450 S/s
1. Anders 021009
2. Anders 021009
yes spec_bias
2002-10-29 13:00:00 -157 2 4 Two-minute bias current sweeps using internal burst at 450 S/s
1. Anders 021009
2. Anders 021009
yes spec_bias
2002-11-01 21:23:00 11820 1 3 By mistake, the failed P1 on SC13 were in E-field mode. This causes the flight s/w to use P12 for calcultating Vsc for HK and IEL, which thus are bad. Science data are OK.
1. Anders 021106
no info
2002-11-02 01:33:00 87060 1 3 By mistake, the failed P1 on SC13 were in E-field mode. This causes the flight s/w to use P12 for calcultating Vsc for HK and IEL, which thus are bad. Science data are OK.
1. Anders 021106
no info
2002-11-03 04:23:00 109080 3 WEC problems probably due to SEUs in DWP.
1. Silvano 021104
2. Keith 021105
no no_tm
2002-11-03 13:34:00 77580 1 WEC problems probably due to SEUs in DWP.
1. Silvano 021104
2. Keith 021105
no no_tm
2002-11-05 08:27:00 20 2 ADCE switchover on SC2.
1. ESOC 021106
2. Keith 021107
no bad_tm
2002-11-05 08:27:20 820 2 Uncontrolled manoeuvre
1. Per-Arne 060331
no bad_data
2002-11-12 01:53:00 840 3 Spacecraft anomaly caused scrambled and lost science data.
1. Keith 021115
2. ESOC 021115
no no_tm
2002-11-12 04:24:00 28740 3 Spacecraft anomaly caused scrambled and lost science data.
1. Keith 021115
2. ESOC 021115
no no_tm
2002-11-12 23:56:24 40 2 P3 stuck after a sweep. no no_p3
2002-11-13 02:34:27 21 2 P3 stuck after a sweep. no no_p3
2002-11-13 03:13:58 43 2 P3 stuck after a sweep. no no_p3
2002-11-13 04:32:59 33 2 P3 stuck after a sweep. no no_p3
2002-11-13 05:52:01 24 2 P3 stuck after a sweep. no no_p3
2002-11-15 03:27:00 100080 2 C2 is bad after a spontanous EFW reset.
1. Per-Arne 060405
no bad_tm
2002-11-21 17:01:00 6060 1 S/c SSR anomaly causing data loss
1. Per-Arne 021203
no no_tm
2002-11-21 23:52:00 54960 1 S/c SSR anomaly causing data loss
1. Per-Arne 021203
no no_tm
2002-11-22 02:19:00 290880 1 WEC data lost during this period.
1. Silvano 021125
2. Keith 021205
no no_tm
2002-11-25 10:15:20 545 3 Data gap due to rain dance procedure.
1. ESOC 021126
no no_tm
2002-11-28 16:10:00 16200 1 2 3 4 Partial BM1 loss.
1. Keith 030410
2. Keith 030406
no no_tm
2002-12-10 12:04:00 6420 1 Attempting to get P1 back to life by deploying 16 both s/c cm more. Deployment worked, but P1 remains unfunctional on SC13.
1. Per-Arne 021210
yes bad_tm
2002-12-10 14:30:00 5160 3 Attempting to get P1 back to life by deploying 16 both s/c cm more. Deployment worked, but P1 remains unfunctional on SC13.
1. Per-Arne 021210
yes bad_tm
2002-12-11 02:02:06 2 4 Bad HX packet
1. Per-Arne 060404
no bad_tm
2002-12-16 16:32:00 45960 4 WEC reset, followed by EFW in h/w default mode (bias -40 nA)
1. Keith 021218
no spec_bias
2002-12-21 20:05:12 12474 2 4 Guard 0 V on probes 3 and 4 yes spec_bias
2002-12-24 14:05:00 16200 1 2 3 4 Partial BM1 loss.
1. Keith 030410
2. Keith 030406
no no_tm
2003-01-06 08:01:56 104416 3 EFW in NULL mode, i.e. no telemetry, due to spurious command.
1. Per-Arne 030107
2. Per-Arne 030108
no no_tm
2003-01-15 12:00:00 300 2 Attitude manoeuvre yes bad_data
2003-01-15 13:00:00 300 4 Attitude manoeuvre yes bad_data
2003-01-15 17:00:00 300 1 Attitude manoeuvre yes bad_data
2003-01-15 18:00:00 300 3 Attitude manoeuvre yes bad_data
2003-01-23 08:13:07 65644 3 WEC off due to DWP problem.
1. Per-Arne 030130
2. Keith 030130
no no_tm
2003-01-24 12:50:36 3606 3 Uplink problem causing setup delay.
1. ESOC 030124
2. Keith 030127
no TBD
2003-01-24 18:54:00 -157 4 WEC switchoff, probably SEU in DWP
1. ESOC 030125
2. Keith 030127
no no_tm
2003-01-30 13:46:00 4860 2 DWP problem? no bad_tm
2003-02-01 11:14:06 -157 3 Partial loss of BM3 dump.
1. Keith 030203
no TBD
2003-02-02 20:00:00 85320 2 4 Guard and puck variations.
1. Anders 030325
yes no_p34
2003-02-03 20:30:00 115500 2 4 Guard and puck variations.
1. Anders 030325
yes no_p34
2003-02-05 22:06:00 27060 1 WEC data loss.
1. ESOC 030211
2. Keith 030210
no no_tm
2003-02-12 11:02:00 10740 1 2 3 4 BM1 data chopped up.
1. Keith 030212
2. ESOC 030212
no no_tm
2003-02-20 11:06:00 1200 4 Data loss due to ground station problem.
1. ESOC 030227
no no_tm
2003-02-24 07:51:00 14340 4 3 x 450 samples/s: E-field V12M, Langmuir mode V3V4M yes spec_bias
2003-02-24 07:51:00 14400 2 Data chopped up. no bad_tm
2003-02-25 15:37:39 2 4 Bad HX telemetry packet. no bad_tm
2003-02-26 10:11:30 600 1 Very bad data because of high density and too large bias current. no info
2003-03-07 23:30:56 1200 3 Eclipse bias +16 nA was commanded 20 minutes too early. yes spec_bias
2003-03-13 07:56:00 -157 2 SC anomaly during upload.
1. ESOC 030314
no TBD
2003-03-14 07:53:14 3 3 Single bad EFW TM packet.
1. Per-Arne 060412
no no_tm
2003-03-19 04:02:00 59760 1 EFW in NULL sampling mode, possibly due to spurious command.
1. Per-Arne 030322
2. Per-Arne 060412
no no_tm
2003-03-19 15:31:36 32258 2 Sync loss between EFW and DWP
1. Keith 030325
no no_tm
2003-03-21 13:07:47 169393 1 EFW switched off by DWP watchdog.
1. Keith 030324
no no_tm
2003-03-22 06:20:00 14520 4 Wrong bias current (+16 nA) due to erroneous eclipse commanding by JSOC.
1. Anders 030409
2. JSOC 030410
no spec_bias
2003-03-22 19:32:00 54396 1 3 Testing 16 bit sweeps and new WEC sequences. Test should have run on SC13, but note problems on SC1 on line above.
1. Anders 030307
yes TBD
2003-03-22 23:59:00 38520 4 EFW in HXONLY mode (by spurious command?). In addition, the HX data for 23:59 - 02:06 are bad because of a bad burst dump. no TBD
2003-03-24 23:01:27 85221 2 P1 first stuck at +70 V, then bad guard voltage causing strange data from 00:07:01 until reset.
1. Anders 060412
no no_p1
2003-03-27 03:54:36 90084 1 3 Testing flight software 2.4 and new WEC sequences (NM1/TM0 and BM1/TM1).
1. Anders 030307
1. Per-Arne 060411
yes TBD
2003-03-27 05:12:00 14400 1 2 3 4 WEC macro problem caused loss of much BM1 data.
1. Keith 030406
no bad_tm
2003-04-05 02:50:00 48420 3 EFW in HXONLY mode (by spurious command?). In addition, HX data are bad 02:50-04:57 because of a bad burst dump. no TBD
2003-04-08 01:30:00 90060 1 3 Testing flight software 2.4 and new WEC sequences (NM1/TM0, BM1/TM1 and BM1/TM2). We have TM2 (22 kbit/s) for the last ten minutes of the BM1 interval. Functions used on each spacecraft in the different sampling mode: SC1: TM0: V23V34L (0x9E) TM1: V23V34M (0x9E) TM2: V2V3V4M (0x1E) SC3: TM0: V2V34L (0x94) TM1: V2V34M (0x9F) TM2: V2V3V34M (0x1F)
1. Per-Arne 060411
yes TBD
2003-04-10 06:26:00 180 1 Very bad data due to high density. no info
2003-04-10 06:49:00 120 1 Very bad data due to high density. no info
2003-04-25 07:00:00 300 2 Attitude manoeuvre. yes bad_data
2003-04-25 08:30:00 300 4 Attitude manoeuvre. yes bad_data
2003-04-25 10:00:00 300 1 Attitude manoeuvre. yes bad_data
2003-04-25 12:00:00 300 3 Attitude manoeuvre. yes bad_data
2003-04-27 09:41:00 16200 1 2 3 4 DWP buffer overflow. no bad_tm
2003-05-03 10:45:00 67440 3 P3 stuck at -10 V.
1. Per-Arne 060402
no bad_tm
2003-05-05 03:57:00 30300 3 WEC switchoff.
1. Keith 030505
2. ESOC 030505
3. Per-Arne 060402
no no_tm
2003-06-04 03:25:54 2086 3 Radial delta-v manoeuvre. yes bad_data
2003-06-06 09:25:00 7200 1 2 3 4 DWP buffer overflow during BM1. no bad_tm
2003-06-09 04:00:00 130 4 Attitude manoeuvre. yes bad_data
2003-06-09 04:30:00 150 2 Attitude manoeuvre. yes bad_data
2003-06-09 06:00:00 170 3 Attitude manoeuvre. yes bad_data
2003-06-11 03:35:00 13920 3 Data gap due to ground station problem.
1. ESOC 030617
2. ESOC 030704
no no_tm
2003-06-18 12:58:22 598664 1 2 3 4 Last expected occurrence of small WEC TM overflow problem that has occurred now and then lately.
1. Keith 030620
no TBD
2003-06-21 06:30:00 140 3 Radial delta-v manoeuvre. yes bad_data
2003-06-21 09:16:00 60 2 Radial delta-v manoeuvre. yes bad_data
2003-06-23 14:01:00 37380 4 All probes saturated at -68 V
1. Per-Arne 030721
no no_tm
2003-06-24 19:35:00 36180 4 Bad data (presumably DWP error)
1. Per-Arne 030721
no no_tm
2003-06-29 02:31:30 20 2 Axial delta-v manoeuvre. yes bad_data
2003-06-29 02:33:30 60 4 Axial delta-v manoeuvre. yes bad_data
2003-07-01 12:39:29 125431 1 HXONLY
1. Per-Arne 030721
no hxonly
2003-07-07 23:41:52 7620 4 Something strange happens to p1, it jumps by about +2 V. no no_p1
2003-07-12 11:27:00 780 1 2 3 4 Sweep tests, bad bias -424 nA on p3,4. Intermediate sweeps on P34, internal burst. (Original request in ref 1 for slow sweeps at perigee was for a time outside data acquisition and had to be replaced, see ref 2)
1. Anders 030623
2. Anders 030626
3. Anders 060418
yes bad_tm
2003-07-12 19:00:00 180 1 2 3 4 Sweep tests, bad bias -424 nA on p3,4. Intermediate sweeps on P34, internal burst.
1. Anders 030623
2. Anders 030626
3. Anders 060418
yes bad_tm
2003-07-13 08:26:00 56160 2 Bad data after spontaneous reset.
1. ESOC 030714
1. Per-Arne 060415
no bad_tm
2003-07-16 14:01:00 3360 1 Bad probe 3 stuck at -68 V. no TBD
2003-07-17 12:30:00 320 1 2 3 4 Slow sweep tests on P34.
1. Anders 030623
2. Anders 030626
3. Anders 060418
yes bad_tm
2003-07-21 16:50:00 360 1 2 3 4 Slow sweep tests.
1. Anders 060418
yes bad_tm
2003-07-23 21:30:39 9 4 Bad sweep handling. The EFW telemetry has bad packets from DWP such that the FDM is not properly decoded for all packets. In this case one FDM containing a sweep=on flag is missing so the sweep data are not removed as they should. no bad_tm
2003-07-24 01:16:00 53040 2 4 Perigee almost-floating bias test on P34: bias -8 nA, guard 0 V. yes spec_bias
2003-07-25 01:41:12 2 3 Bad HX packet. no bad_tm
2003-07-25 05:07:00 10200 4 WEC off
1. Per-Arne 030725
2. ESOC 030725
no no_tm
2003-07-29 07:00:00 9720 2 EFW bad data
1. Per-Arne 030730
no bad_tm
2003-08-03 03:16:00 114060 1 WEC off
1. ESOC 030804
2. ESOC 030804
no no_tm
2003-08-14 14:12:00 33960 1 EFW data gaps
1. Per-Arne 030819
no bad_tm
2003-08-20 07:50:11 45600 2 P1 stuck at +68 V until 09:42:23, then bad guard until reset at 20:30.
1. Per-Arne 060415
no no_p1
2003-08-26 17:11:29 9916 2 P3 goes positive after a sweep. no no_p3
2003-09-26 18:15:00 290 2 Attitude manoeuvre. yes bad_data
2003-09-26 19:10:00 290 3 Attitude manoeuvre. yes bad_data
2003-10-07 06:00:00 2400 2 4 Slow sweeps on P34
1. Anders 030918
yes spec_bias
2003-10-08 15:06:00 2220 4 WEC power cycling for WHISPER problem.
1. ESOC 031008
no no_tm
2003-10-09 04:52:00 41340 4 WEC data gap due to DWP problem?
1. Per-Arne 031015
2. Keith 031015
no no_tm
2003-10-10 17:09:00 2280 2 WEC power cycling for DWP latchup, curing a STAFF cal problem
1. Keith 031010
2. ESOC 031010
no no_tm
2003-10-13 22:15:18 2 3 Single bad HX packet. no bad_tm
2003-10-15 01:48:11 121795 2 WEC emergency mode due to DWP problem, causing 0 bias.
1. Keith 031016
no spec_bias
2003-10-16 11:38:06 2071 2 WEC power cycle to clear up DWP problem.
1. ESOC 031016
2. Keith 031016
yes no_tm
2003-10-21 15:29:00 71520 4 NULL mode. no no_tm
2003-10-27 01:56:05 2 3 Single bad HX packet. no bad_tm
2003-10-28 15:21:00 120 3 WEC problems due to proton event
1. ESOC 031029
2. ESOC 031029
no no_tm
2003-10-28 16:17:00 22560 4 WEC problems due to proton event
1. ESOC 031029
2. ESOC 031029
no no_tm
2003-10-29 01:58:00 46260 1 All Cluster data lost due to proton event impact on s/c systems
1. ESOC 031030
2. AR-051 (pdf)
3. ESOC 031029
no no_tm
2003-10-29 08:30:00 30720 3 WEC and EFW problems during solar proton event
1. ESOC 030129
2. Per-Arne 031029
3. ESOC 031029
no bad_tm
2003-10-29 16:57:00 10320 3 No V32 available as FSW 2.4 appears not to have loaded properly.
1. Yuri 060424
no TBD
2003-10-30 17:27:00 1680 2 WEC power cycling following solar proton event problems
1. ESOC 031030
2. ESOC 031029
no no_tm
2003-11-12 13:57:00 37440 1 Bad data after spontaneous reset. no bad_data
2003-11-27 10:26:00 30960 1 Probe 3 stuck at around -10 V (just like probe 1). It is interesting that the problem went away at the first Whisper sounding in BM (the Whisper soundings in NM did not help)! Both p3, p34 and p32 are affected. no no_p3
2003-12-11 20:55:00 10800 1 2 3 4 Buffer overflow in DWP during BM. no bad_tm
2003-12-13 04:50:00 62640 3 Bad data after spontaneous EFW reset. no bad_data
2003-12-13 18:43:00 12600 2 Bias test, P34 guard zero
1. Anders 031125
2. Anders 031125
yes spec_bias
2003-12-13 19:10:00 7380 1 2 3 4 Sweep interval 10 minutes
1. Anders 031125
2. Anders 031125
yes TBD
2003-12-14 00:08:00 220620 3 WBD problem turned WEC off. Several tests followed up to Dec 18.
1. ESOC 031214
2. Keith 031215
3. ESOC 031215
4. Keith 031215
5. ESOC 031216
6. Keith 031219
7. Jolene 040129
no no_tm
2003-12-14 07:30:00 56520 2 3 4 Sweep interval 30 minutes. Should have run also on SC3, but the WBD problem stopped this.
1. Anders 031125
2. Anders 031125
yes TBD
2003-12-16 03:30:00 360 3 Slow sweeps on P2. Should have run also on SC3, but the WBD problem stopped this. yes spec_bias
2003-12-21 06:52:34 2 3 Single bad HX packet. no bad_tm
2004-01-07 15:00:00 300 2 Attitude manoeuvre
1. MOR 179 (PDF)
yes bad_data
2004-01-07 15:30:00 300 3 Attitude manoeuvre
1. MOR 179 (PDF)
yes bad_data
2004-01-07 17:30:00 300 1 Attitude manoeuvre
1. MOR 179 (PDF)
yes bad_data
2004-01-07 18:15:00 300 4 Attitude manoeuvre
1. MOR 179 (PDF)
yes bad_data
2004-01-08 03:05:00 82085 1 2 4 DWP/WBD command problem
1. Keith 040109
2. ESOC 040109
no no_tm
2004-01-08 03:05:00 37170 3 DWP/WBD command problem
1. Keith 040109
2. ESOC 040109
no no_tm
2004-01-09 01:42:00 665 3 DWP/WBD command problem
1. Keith 040109
2. ESOC 040109
no no_tm
2004-01-29 22:06:00 571140 1 EFW problem
1. Per-Arne 040205
2. Keith 040205
3. Per-Arne 040205
4. ESOC 040205
no bad_tm
2004-02-05 13:23:00 79860 3 NULL sampling mode no bad_tm
2004-02-15 21:27:00 20940 3 HXONLY sampling mode
1. Per-Arne 060427
no hxonly
2004-02-26 07:32:00 197940 1 Bad EFW data. It is not clear what happened to EFW, but all quantities are stuck at fixed values. no bad_tm
2004-02-27 14:00:00 72780 3 WEC off by onboard monitoring
1. Keith 040228
no no_tm
2004-03-14 08:18:00 112800 1 EFW/DWP communication problem
1. Per-Arne 040315
2. Keith 040315
3. ESOC 040315
4. Keith 040315
no no_tm
2004-03-26 16:21:00 77940 1 Bad data after spontaneous EFW reset no bad_data
2004-03-29 00:24:12 2 3 Single bad HX packet no bad_tm
2004-04-05 22:34:00 9060 2 Probe 1 at large positive potential but not quite stuck no no_p1
2004-04-08 06:58:32 960 4 Probe 1 suddenly goes positive (saturated) and then comes back but still stays bad at positive values until it is restored by a bias sweep. no no_p1
2004-04-11 00:45:26 876 2 Whisper continuously active no bad_data
2004-04-12 20:40:05 2 3 Single bad HX packet no bad_tm
2004-04-23 10:39:29 2588 4 WEC power cycle to clear up DWP problem.
1. ESOC 040503
no no_tm
2004-05-02 03:30:00 5340 1 Radial delta-v manoeuvre yes bad_data
2004-05-02 03:48:00 4740 3 Radial delta-v manoeuvre yes bad_data
2004-05-04 06:20:00 254400 2 Shadow pulses in data due to Solar Aspect Angle (SAA) less than 93.2 degrees no bad_data
2004-05-04 06:20:00 276000 4 Shadow pulses in data due to Solar Aspect Angle (SAA) less than 93.2 degrees no bad_data
2004-05-07 05:00:00 320 2 Attitude manoeuvre yes bad_data
2004-05-07 06:00:00 305 3 Attitude manoeuvre yes bad_data
2004-05-07 10:00:00 340 1 Attitude manoeuvre yes bad_data
2004-05-07 11:00:00 355 4 Attitude manoeuvre yes bad_data
2004-05-14 17:18:00 54360 2 WEC switched off by s/c
1. ESOC 040515
no no_tm
2004-05-18 10:00:00 15600 1 2 4 Very bad data due to high density no info
2004-05-19 06:00:00 30 2 Attitude manoeuvre yes bad_data
2004-05-19 07:00:00 40 3 Attitude manoeuvre yes bad_data
2004-05-19 09:00:00 45 1 Attitude manoeuvre yes bad_data
2004-05-19 10:00:00 70 4 Attitude manoeuvre yes bad_data
2004-05-19 13:49:16 155001 4 WEC switched off by s/c
1. ESOC 040527
no no_tm
2004-05-20 01:45:00 30120 1 Bad data. It looks like we have a lot of digital bit errors in the telemetry, which causes spikes in the measured data but also timing errors. It is probably an internal DWP problem, since it is cleared by the "stop macro" command at 10:07. no bad_tm
2004-05-21 05:00:00 60 2 Spin-down manoeuvre yes bad_data
2004-05-21 06:00:00 20 3 Spin-down manoeuvre yes bad_data
2004-05-21 08:30:00 20 1 Spin-up manoeuvre yes bad_data
2004-05-21 09:30:00 20 4 Spin-down manoeuvre yes bad_data
2004-05-21 09:34:00 37020 1 Probe 1 at large positive potential but not quite stuck, problem was cleared by EFW reset no no_p1
2004-05-29 09:35:00 10800 1 2 3 4 Bad WEC data due to DWP buffer overflow in BM no bad_tm
2004-05-30 03:46:30 180 1 Radial delta-v manoeuvre yes bad_data
2004-06-05 20:59:00 124200 4 Bad data after spontaneous EFW reset no bad_data
2004-06-07 09:16:55 10 3 Axial trim manoeuvre yes bad_data
2004-06-11 01:05:30 19500 2 WEC off
1. ESOC 040611
2. Keith 040611
no no_tm
2004-06-13 18:17:19 12971 3 Probe 3 temporarily broken, for reasons to be investigated.
1. Per-Arne 060309
no no_p3
2004-06-13 21:53:30 70950 1 S/C problem: CTU switchover
1. ESOC 040614
2. Keith 040614
3. ESOC 040614
4. Keith 040615
5. ESOC 040615
6. ESOC 040616
7. Keith 040616
8. ESOC 040616
9. AR-054 (pdf)
no no_tm
2004-06-14 00:16:00 47940 2 Very bad data due to high density no info
2004-06-14 00:19:00 43980 4 Very bad data due to high density no info
2004-06-14 02:37:00 18360 3 Very bad data due to high density no info
2004-06-16 09:57:00 3600 1 2 3 4 Lunar eclipse
1. ESOC report (pdf)
yes TBD
2004-06-21 14:01:36 3 4 Single bad HX packet no bad_tm
2004-07-03 03:14:00 2340 2 Data loss due to s/c problem
1. AR-057 (pdf)
no no_tm
2004-07-09 10:35:00 56400 2 HXONLY sampling mode
1. Per-Arne 040709
2. Per-Arne 040710
no hxonly
2004-07-19 22:59:44 3 4 Single bad HX packet no bad_tm
2004-07-21 01:07:00 26760 4 NULL sampling mode no bad_tm
2004-07-21 19:20:04 3 1 Single bad TM packet no bad_tm
2004-07-26 05:12:00 8280 2 Probe 1 at large positive potential but not quite stuck, problem was cleared by sweep no no_p1
2004-07-28 11:19:00 9060 2 Probe 1 at large positive potential but not quite stuck, problem was cleared by sweep no no_p1
2004-08-09 10:05:00 6600 1 2 3 4 Bad data due to DWP buffer overflow in BM no bad_tm
2004-08-14 04:10:00 7200 1 2 3 4 Bad data due to DWP buffer overflow in BM no bad_tm
2004-08-14 17:44:43 3 3 Single bad TM packet no bad_tm
2004-08-14 23:00:00 69180 3 NULL sampling mode
1. Per-Arne 040816
2. Per-Arne 040816
no bad_tm
2004-08-18 12:48:00 68640 3 WEC switched off by s/c. Note: Switch-off time in first ESOC mail is wrong.
1. ESOC 040819
2. ESOC 040825
no no_tm
2004-08-19 08:06:00 18840 3 WEC switched off by s/c. Note: Switch-off time in first ESOC mail is wrong.
1. ESOC 040819
2. ESOC 040825
no no_tm
2004-08-21 12:35:00 51600 2 WEC switched off by s/c
1. ESOC 040823
2. Keith 040831
no no_tm
2004-08-22 00:15:00 270 1 Attitude manoeuvre yes bad_data
2004-08-22 01:00:00 270 4 Attitude manoeuvre yes bad_data
2004-08-22 02:30:00 260 3 Attitude manoeuvre yes bad_data
2004-08-31 19:55:00 10800 1 2 3 4 Bad data in BM due to DWP buffer overflow no bad_tm
2004-09-05 08:33:00 68220 1 2 3 4 Payload off during long (apogee) eclipse yes no_tm
2004-09-07 16:23:00 72630 1 2 3 4 Payload off during long (apogee) eclipse yes no_tm
2004-09-08 21:52:33 2 4 Single bad TM packet no bad_tm
2004-09-10 01:00:50 71660 1 2 3 4 Payload off during long (apogee) eclipse yes no_tm
2004-09-16 19:54:08 183078 1 EFW data lost for almost 51 h, presumably DWP problem.
1. ESOC 040923
2. Anders 060413
3. Anders 060508
no no_tm
2004-09-16 21:20:00 10800 2 3 4 BM1 data chopped up, presumably DWP problem.
1. ESOC 040923
2. Anders 060413
3. Anders 060508
no bad_tm
2004-09-17 16:29:40 2 3 Single bad TM packet no bad_tm
2004-09-18 16:10:45 34225 2 Data lost due to s/c problem
1. ESOC 040920
2. AR-061 (pdf)
3. ESOC 040923
no no_tm
2004-09-18 22:45:26 1758 1 FSW 2.1 run instead of 2.4 no TBD
2004-09-18 23:14:44 121242 1 EFW data lost or bad, presumably DWP problem
1. ESOC 040923
2. Anders 060413
3. Anders 060508
no no_tm
2004-09-19 22:14:00 39960 3 NULL sampling mode
1. ESOC 040923
2. Anders 060413
3. Anders 060508
no bad_tm
2004-09-20 09:17:08 112697 1 WEC switched off by s/c
1. ESOC 040921
2. ESOC 040921
3. Keith 040921
4. Per-Arne 040922
5. ESOC 040923
6. Anders 060413
7. Anders 060508
no no_tm
2004-09-21 19:04:00 150120 4 DWP problem, cured by WEC power cycling
1. ESOC 040922
2. Per-Arne 040922
3. ESOC 040930
no bad_tm
2004-09-24 13:48:00 48060 1 Bad data after spontaneous EFW reset no bad_data
2004-09-25 20:23:32 32308 3 WEC off
1. ESOC 040930
no no_tm
2004-09-27 04:32:50 21970 3 NULL sampling mode no bad_tm
2004-10-05 17:27:10 2 1 Single bad TM packet no bad_tm
2004-10-11 11:08:44 2 3 Single bad TM packet no bad_tm
2004-10-13 11:13:50 28150 3 HXONLY sampling mode no hxonly
2004-10-18 17:00:00 178620 4 S/c problem (CTU switchover)
1. ESOC 041020
2. AR-062 (pdf)
no no_tm
2004-10-20 10:56:43 2 1 Single bad TM packet no bad_tm
2004-10-26 16:19:16 2 3 Single bad TM packet no bad_tm
2004-10-30 19:15:00 69430 4 HXONLY sampling mode no hxonly
2004-10-31 01:45:10 219590 3 S/c problem (CTU switchover)
1. ESOC 041101
2. AR-063 (PDF)
3. ESOC 041102
no no_tm
2004-11-02 00:43:00 140760 4 S/c problem (SSR controller failure)
1. ESOC 041104
2. AR-064 (PDF)
no no_tm
2004-11-06 15:10:00 171000 3 WEC off following spurious WBD command.
1. ESOC 041108
2. Keith 041108
3. ESOC 041108
no no_tm
2004-11-07 10:00:00 20220 1 2 4 Very bad data due to high density no info
2004-11-09 08:22:49 110251 1 Bad EFW TM followed by WEC off by monitoring.
1. ESOC 041118
no bad_tm
2004-11-10 23:22:00 28380 1 Bad data after spontaneous EFW reset no bad_data
2004-11-12 16:26:19 106741 1 DWP time problem gave bad EFW TM, followed by WEC off by monitoring.
1. ESOC 041118
no bad_tm
2004-11-27 23:53:00 1 1 Burst dump lost due to dump made in BM1 rather than BM3.
1. Keith 041129
2. ESOC 041130
no no_bm3
2004-11-28 14:18:00 19860 1 2 4 Very bad data due to high density no info
2004-11-28 15:34:00 13800 3 Very bad data due to high density no info
2004-11-30 09:09:00 103800 3 WEC emergency mode following DWP reset; -40 nA on all probes.
1. ESOC 041209
no bad_tm
2004-12-01 13:59:00 11160 3 DWP problem gave bad EFW TM, followed by WEC off by monitoring.
1. ESOC 041209
no bad_tm
2004-12-01 17:19:00 14760 3 WEC off by monitoring
1. ESOC 041209
no no_tm
2004-12-05 05:12:00 9300 1 2 4 Very bad data due to high density no info
2004-12-12 23:03:00 2400 1 2 3 4 Very bad data due to high density no info
2004-12-24 14:00:00 94680 2 WEC switched off by onboard monitoring
1. Keith 041231
no no_tm
2004-12-31 04:06:00 81370 3 NULL sampling mode no bad_tm
2005-01-08 07:59:00 5940 1 2 3 4 Very bad data due to high density no info
2005-01-17 17:33:00 46800 4 Bad data after spontaneous EFW reset no bad_data
2005-01-17 18:27:00 43200 3 NULL sampling mode no bad_tm
2005-01-20 07:11:00 29820 3 No EFW data
1. Per-Arne 050121
2. ESOC 050121
no no_tm
2005-01-20 15:28:00 73920 3 Bad data, all probes latched at -68 V
1. Per-Arne 050121
2. ESOC 050121
no bad_data
2005-01-22 21:35:00 11160 3 Bad WEC telemetry
1. Per-Arne 050121
2. ESOC 050121
no bad_tm
2005-01-27 20:35:00 13980 1 2 3 4 Very bad data due to high density no info
2005-02-05 16:59:00 16080 1 2 3 4 Very bad data due to high density no info
2005-02-17 23:59:00 660 1 2 4 Very bad data due to high density no info
2005-02-19 20:08:29 3 2 Single bad HX packet no bad_tm
2005-02-22 08:57:00 4440 1 2 3 4 Very bad data due to high density no info
2005-02-23 06:00:00 1260 1 2 4 Very bad data due to high density no info
2005-03-04 02:20:00 16440 4 Bad data after spontaneous EFW reset no bad_data
2005-03-06 06:10:00 9300 1 2 4 Very bad data due to high density no info
2005-03-10 07:28:57 2 3 Single bad HX packet no bad_tm
2005-03-13 06:55:00 22020 1 2 3 4 Intermittently very bad data due to high density no info
2005-03-14 01:10:00 12660 1 2 3 4 Very bad data due to high density no info
2005-03-14 10:30:00 260 1 Attitude manoeuvre yes bad_data
2005-03-14 11:00:00 260 4 Attitude manoeuvre yes bad_data
2005-03-14 15:30:00 250 2 Attitude manoeuvre yes bad_data
2005-03-14 16:00:00 250 3 Attitude manoeuvre (wrong time given in Mission Operations Report) yes bad_data
2005-03-20 09:28:00 46620 4 Bad data after spontaneous EFW reset no bad_data
2005-03-23 13:08:00 9780 1 Very bad data due to high density no info
2005-03-23 13:52:00 7260 2 Very bad data due to high density no info
2005-03-23 14:06:00 6780 3 Very bad data due to high density no info
2005-03-25 21:14:53 8 3 Bad TM due to unknown reasons no bad_tm
2005-04-03 17:10:00 18720 1 2 3 4 Very bad data due to high density no info
2005-04-04 01:57:00 9300 3 Apparent DWP buffer overflow no bad_tm
2005-04-05 11:55:00 109920 3 Bad data after spontaneous EFW reset no bad_data
2005-04-06 17:17:15 3 2 Two bad HX packets no bad_tm
2005-04-10 06:47:00 107400 1 3 P2 at 25 S/s rather than P23, for SC13 only yes no_p32
2005-04-11 12:38:00 15060 1 2 3 4 Very bad data due to high density no info
2005-04-16 08:50:00 9000 1 2 3 4 Bad data due to DWP buffer overflow in BM no bad_tm
2005-04-21 12:41:00 2640 2 WEC switched off by onboard monitoring
1. ESOC 050421
no no_tm
2005-04-29 19:31:00 8760 1 2 3 4 Very bad data due to high density no info
2005-05-05 10:00:00 10800 1 2 3 4 Bad data due to DWP buffer overflow in BM no bad_tm
2005-05-07 18:28:00 11400 1 2 3 4 Very bad data due to high density no info
2005-05-10 11:05:00 5400 1 2 3 4 Bad data due to DWP buffer overflow in BM no bad_tm
2005-05-11 22:43:40 2 3 Single bad HX packet no bad_tm
2005-05-17 01:08:00 7500 2 Probe 1 at large positive potential but not quite stuck, problem was cleared by sweep no no_p1
2005-05-17 01:14:00 7140 4 Probe 1 at large positive potential but not quite stuck, problem was cleared by sweep no no_p1
2005-05-17 16:16:00 51060 2 WEC switched off by onboard monitoring
1. ESOC 050518
no no_tm
2005-05-23 10:33:00 22260 3 HXONLY sampling mode no hxonly
2005-05-23 16:44:00 19740 3 HXONLY sampling mode and bad data due to low density saturation no bad_data
2005-05-23 22:13:00 21600 3 HXONLY sampling mode no hxonly
2005-05-27 01:42:00 123780 1 2 3 4 Very bad data due to high density no info
2005-06-01 03:59:00 2880 4 Axial manoeuvre no bad_data
2005-06-01 07:46:00 2940 3 Axial manoeuvre no bad_data
2005-06-08 04:16:00 7680 4 Radial manoeuvre no bad_data
2005-06-11 15:50:00 87600 3 Apparent DWP buffer overflow no bad_tm
2005-06-12 07:26:00 31560 1 Probe 1 incorrectly set to E-field mode with zero bias (used wrong sequence SWECJ305 after interval of no data taking) and stuck at -68 V no no_p1
2005-06-12 07:47:00 34140 2 Very bad data due to high density no info
2005-06-14 00:35:00 368700 2 Shadow pulses in data due to Solar Aspect Angle (SAA) less than 93.2 degrees yes bad_data
2005-06-15 09:08:00 9000 4 WEC overflow due to wrong correlator mode
1. ESOC 050617
no bad_tm
2005-06-17 07:42:00 9360 3 No EFW data due to DWP problem? no no_tm
2005-06-18 07:00:00 60 2 Attitude manoeuvre yes bad_data
2005-06-18 08:00:00 300 3 Attitude manoeuvre yes bad_data
2005-06-20 06:00:00 20 1 Spin-up manoeuvre yes bad_data
2005-06-20 09:00:00 30 4 Spin-up manoeuvre yes bad_data
2005-06-22 18:18:00 63720 1 2 3 4 Very bad data due to high density no info
2005-06-24 20:24:00 1080 2 Axial manoeuvre yes bad_data
2005-06-29 14:56:30 30 2 Axial manoeuvre yes bad_data
2005-06-30 07:01:35 5 1 Radial manoeuvre yes bad_data
2005-06-30 09:46:20 90 4 Radial manoeuvre yes bad_data
2005-07-06 18:10:00 10 2 Axial manoeuvre yes bad_data
2005-07-06 21:35:00 30 3 Axial manoeuvre yes bad_data
2005-07-10 11:12:00 2760 1 2 3 4 Very bad data due to high density no info
2005-07-24 00:22:00 71520 4 Bad data after spontaneous EFW reset. no bad_data
2005-07-24 20:14:00 17940 4 Data loss due to SC CTU switchover at 2005-07-26 07:01.
1. ESOC 050726
2. ESOC 050726
3. Per-Arne 050726
4. AR-70 (pdf)
no no_tm
2005-07-25 01:13:00 6300 4 Bad data after spontaneous EFW reset. no bad_data
2005-07-25 02:58:00 93360 4 Data loss due to SC CTU switchover at 2005-07-26 07:01.
1. ESOC 050726
2. ESOC 050726
3. Per-Arne 050726
4. AR-70 (pdf)
no no_tm
2005-07-26 07:01:00 27600 4 Data loss due to SC CTU switchover at 2005-07-26 07:01.
1. ESOC 050726
2. ESOC 050726
3. Per-Arne 050726
4. AR-70 (pdf)
no no_tm
2005-08-10 12:42:00 16380 4 S/c test (CTU-B)
1. ESOC 050805
2. ESOC 050811
yes no_tm
2005-08-10 18:30:00 1200 3 WEC power cycle follwoing DWP latchup.
1. Keith 050805
yes no_tm
2005-08-19 16:45:00 46980 4 Bad data after spontaneous EFW reset no bad_data
2005-09-07 15:01:44 20683 4 No HK data due to s/c problem. No science data due to (planned) eclipse ops
1. ESOC 050921
2. AR-73
no no_tm
2005-09-14 13:04:00 4140 3 Bias = +4 nA in penumbra eclipse yes bad_data
2005-09-14 13:17:00 4080 4 Bias = +4 nA in penumbra eclipse yes bad_data
2005-09-14 17:55:18 136050 3 WEC off following spurious WBD command
1. Keith 050916
no no_tm
2005-09-16 01:35:54 7920 4 Data overwritten onboard
1. ESOC 050920
no no_tm
2005-09-16 13:10:00 12240 4 Probe 2 stuck at -68 V, problem was cleared by sweep no no_p2
2005-09-27 20:47:00 3360 2 Very bad data due to high density no info
2005-09-27 21:00:00 5580 1 Very bad data due to high density no info
2005-09-27 23:48:00 3660 3 Very bad data due to high density no info
2005-09-28 00:02:00 3480 4 Very bad data due to high density no info
2005-10-01 03:31:00 151140 2 WEC problem, switched off by onboard monitoring at 2005-10-01 05:49
1. ESOC 051003
no bad_data
2005-10-05 19:00:00 300 2 Attitude manoeuvre. yes bad_data
2005-10-05 20:00:00 300 4 Attitude manoeuvre. yes bad_data
2005-10-05 23:30:00 300 1 Attitude manoeuvre. yes bad_data
2005-10-06 00:30:00 320 3 Attitude manoeuvre. yes bad_data
2005-10-07 08:27:00 11880 3 Very bad data due to high density. no info
2005-10-07 08:27:00 14040 4 Very bad data due to high density. no info
2005-10-07 09:18:00 5820 2 Very bad data due to high density. no info
2005-10-07 09:19:00 5520 1 Very bad data due to high density. no info
2005-10-31 02:58:00 31980 2 DWP buffer overflow
1. ESOC 051031
no TBD
2005-10-31 13:38:00 163800 4 DWP latchup, eventually cleared up by WEC power cycle
1. ESOC 051102
no TBD
2005-11-09 05:36:00 25080 1 Apparent DWP buffer overflow no bad_tm
2005-11-11 00:20:00 11400 1 2 3 4 Very bad data due to high density no info
2005-11-18 05:50:00 70500 2 Very bad data due to high density no info
2005-11-18 06:30:00 69900 3 Very bad data due to high density no info
2005-11-18 06:30:00 81900 4 Very bad data due to high density no info
2005-11-18 09:55:00 57600 1 Very bad data due to high density no info
2005-11-19 10:15:52 10 2 Phasing manoeuvre yes bad_data
2005-11-28 09:03:00 37440 2 Very bad data due to high density no info
2005-11-28 09:03:00 39480 1 3 Very bad data due to high density no info
2005-11-28 09:03:00 43020 4 Very bad data due to high density no info
2005-11-29 13:26:00 17160 2 Bad data after spontaneous EFW reset, cured by WEC power cycling
1. Keith 051129
no bad_data
2005-11-29 23:54:00 16320 1 Bad data after spontaneous EFW reset no bad_data
2005-11-30 14:38:00 91920 1 DWP problem followed by WEC off by onboard monitoring
1. ESOC 051201
no bad_data
2005-12-07 11:58:00 191820 1 Instrument hangup
1. Per-Arne 051209
2. Per-Arne 051213
3. Keith 051214
no bad_tm
2005-12-08 13:17:00 7560 4 Probe 1 positive, starts slightly positive and gradually increases to +68 V, problem cleared by sweep no no_p1
2005-12-09 06:22:00 94080 2 3 4 Very bad data due to high density no info
2005-12-09 17:15:00 54900 1 Very bad data due to high density no info
2005-12-13 05:25:00 1500 1 3 Slow sweeps on P2 SC13 in plasmasphere/plasmapause. Sampled V2 and V34 2005-12-12 01:30 - 2005-12-13 09:32.
1. Anders 051117
2. Anders 051214
yes spec_bias
2005-12-13 07:13:00 8340 3 Bad data after spontaneous EFW reset no bad_data
2005-12-16 02:11:38 2 3 Single bad HX packet no bad_tm
2005-12-16 05:10:00 133800 1 Very bad data due to high density no info
2005-12-16 05:10:00 130680 2 Very bad data due to high density no info
2005-12-16 05:10:00 103020 3 Very bad data due to high density no info
2005-12-16 05:10:00 140700 4 Very bad data due to high density no info
2005-12-17 09:47:00 124260 3 Bad data after spontaneous EFW reset
1. Keith 051219
no bad_data
2005-12-18 11:50:00 18600 2 Very bad data due to high density no info
2005-12-18 14:14:00 18180 1 Very bad data due to high density no info
2005-12-18 14:14:00 18060 4 Very bad data due to high density no info
2005-12-19 00:32:00 14880 1 2 3 4 Very bad data due to high density no info
2005-12-19 10:36:38 2 3 Single bad HX packet no bad_tm
2005-12-23 02:55:22 2 3 Single bad HX packet no bad_tm
2005-12-24 06:40:00 52800 1 2 3 4 Intermittently very bad data due to high density no info
2005-12-26 21:25:00 29100 1 2 3 4 Very bad data due to high density no info
2006-01-06 13:00:00 300 4 Maneuver.
yes bad_data
2006-01-08 07:05:00 23100 4 Unidentified DWP problem caused EFW and WHISPER passive data loss
1. Keith 060117
no no_tm
2006-01-08 07:05:02 23130 4 No EFW TM. DWP or EFW problem, not investigated.
1. ESOC 060112
no no_tm
2006-01-09 13:50:00 2460 4 Testing sweeps at different sweep rates
1. Anders 051214
yes spec_bias
2006-01-10 19:00:00 1800 3 Testing coarse sweeps in plasmasphere
1. Anders 051214
yes spec_bias
2006-01-10 20:00:00 3600 4 Testing fine sweeps in plasmasphere
1. Anders 051214
yes spec_bias
2006-02-10 21:22:18 453 1 SC1 raindance procedure (note other simultaneous problems above).
1. ESOC 060210
yes no_tm
2006-02-11 21:02:00 236717 1 Several, probably connected, issues: (1) EFW data missing from 2006-02-10 21:14:20, probably DWP problem. (2) WEC off due to DWP latchup 2006-02-11 21:02 until manual reconfiguration. The raindance procedure (see below) at 2006-02-10 21:22:18 does no change to the lack of EFW data.
1. Per-Arne 060211
2. ESOC 060213
3. ESOC 060224
no no_tm
2006-02-12 13:10:00 300 4 Garbled telemetry.
no bad_tm
2006-02-12 15:30:00 30600 4 Bias set to -40 and 0. Problems in DSC data.
no bad_bias
2006-02-14 07:12:32 102 4 EFW bad data due to DWP problem (SEU).
1. Keith 060214
no bad_tm
2006-02-14 22:21:06 3 4 One bad EFW packet, due to DWP problem (SEU).
1. Keith 060214
no bad_tm
2006-02-14 22:22:50 60900 4 EFW off due to DWP problem (SEU).
1. Keith 060214
no no_tm
2006-02-24 12:44:45 145 3 WEC off by DWP voltage monitoring.
1. Keith 060224
2. ESOC 060302
no no_tm
2006-02-24 12:47:10 21560 3 WEC emergency mode following DWP reset, -40 nA on all probes.
1. Keith 060224
2. ESOC 060302
no spec_bias
2006-03-04 17:36:00 66660 3 WEC off by DWP voltage monitoring.
1. ESOC 060308
no no_tm
2006-03-08 08:31:56 61084 3 Bad bias settings, possibly due to SEU.
no bad_bias
2006-03-08 20:51:58 11282 1 Bad bias settings.
no bad_bias
2006-03-14 05:44:37 24900 1 WEC off by DWP voltage monitoring.
1. ESOC 060314
no no_tm
2006-03-14 08:41:08 562 4 Bad packet header times. no bad_tm
2006-03-14 12:34:30 1500 4 Bad packet header times. no bad_tm
2006-03-14 21:53:00 88400 4 Payload off following CTU switchover.
1. ESOC 060315
2. ESOC 060315
3. ESOC 060322
no no_tm
2006-03-15 09:34:00 5040 4 RTU switchover recovery plan.
1. ESOC 060322
yes no_tm
2006-03-16 14:18:41 60 3 Battery gas release.
1. Per-Arne 060330
2. Per-Arne 060330
3. ESOC 060330
4. Pat Daly 060330
5. ESOC 060330
6. Per-Arne 060330
7. Iannis 060330
8. Pat Daly 060331
9. Per-Arne 060331
10. Patrick Canu 060407
11. Iannis 060411
no TBD
2006-03-17 08:30:30 96990 2 Garbled telemetry. Data unusable.
no bad_tm
2006-03-19 09:42:10 19070 3 Probe saturation latchup. p2 goes positive, others go negative.
no bad_data
2006-03-20 00:04:08 60 1 Battery gas release.
1. Per-Arne 060330
2. Per-Arne 060330
3. ESOC 060330
4. Pat Daly 060330
5. ESOC 060330
6. Per-Arne 060330
7. Iannis 060330
8. Pat Daly 060331
9. Per-Arne 060331
10. Patrick Canu 060407
11. Iannis 060411
no TBD
2006-03-24 15:38:52 341168 1 SC1 battery problem in short eclipse.
1. ESOC 060327
2. ESOC 060405
3. AR-088
no no_tm
2006-03-30 12:22:26 167974 4 WEC switched off by voltage monitoring.
1. ESOC 060403
2. ESOC 060405
no no_tm
2006-04-02 04:39:30 125370 3 WEC switched off because of an uncommanded WBD on following a DWP watchdog reset. During the off period, a spacecraft SSR controller failure occured and was cured.
1. ESOC 060405
2. AR-089 (pdf)
no no_tm
2006-04-07 06:59:00 360 3 Maneuver.
yes bad_data
2006-04-07 07:59:00 360 2 Maneuver.
yes bad_data
2006-04-07 12:59:00 360 1 Maneuver.
yes bad_data
2006-04-07 20:29:00 360 4 Maneuver.
yes bad_data
2006-04-08 11:48:30 42660 1 Probe 3 latched at -68 V.
no bad_data
2006-04-20 14:33:32 532 2 Problem with the data (looks digital, see below). no bad_tm
2006-04-20 14:34:19 75581 2 DWP problem finally cured by WEC power cycle. Problem may have started earlier than stated here, and also end somewhat earlier (or possibly later).
1. ESOC 060421
no no_tm
2006-04-21 19:09:30 12600 2 Serious problems with the data: digital hiccups as well as bad bias. no bad_data
2006-05-09 07:11:56 17441 4 WEC turned off for s/c (CTU-B) test.
1. ESOC 060509
yes no_tm
2006-05-09 12:10:11 3 4 Single bad EFW data packet.
1. Per-Arne 060514
no bad_tm
2006-05-10 05:01:00 68340 3 Bias set to -40 nA.
no bad_bias
2006-05-13 00:56:48 2 3 Bad packet.
no bad_data
2006-05-24 12:08:20 4140 2 Data gaps following DWP reset, cured by WEC power cycle.
1. ESOC 060601
no bad_tm
2006-06-07 03:21:00 132060 3 EFW digital problems.
1. Per-Arne 060608
no bad_tm
2006-06-08 16:02:00 1680 3 WEC power cycle.
1. Per-Arne 060609
yes no_tm
2006-06-12 10:03:24 84553 4 WEC off by DWP voltage monitoring.
1. ESOC 060613
no no_tm
2006-06-18 05:52:00 96540 4 Bias set to -40 nA.
no bad_bias
2006-06-23 15:44:00 39600 3 Probe latchup. p1,3 go negative, p2,4 go positive.
no bad_data
2006-06-29 16:30:00 19320 1 Probe latchup on p3.
no no_p3
2006-06-30 23:48:00 22920 3 Bad bias settings. (-40 nA)
no bad_bias
2006-07-05 15:08:00 33120 3 No TM -- DWP commanding problem?
1. ESOC 060725
no no_tm
2006-07-07 14:22:00 924810 1 DWP memory latchup
1. Keith 060714
2. ESOC 060718
no TBD
2006-07-09 07:41:00 5760 1 Bad DWP, EFW and STAFF-SA TLM received on SC1 on DOY 190 from 07.41 to 09.17 following watchdog reset.
1. ESOC 060725
no bad_tm
2006-07-12 12:30:00 1200 3 Maneuver. yes bad_data
2006-07-15 05:00:02 2 3 Bad packet. no bad_tm
2006-07-21 06:32:00 3000 1 WEC power cycle following DWP memory latchup.
1. ESOC 060725
no no_tm
2006-07-31 07:09:00 86400 1 WEC off by DWP voltage monitoring.
1. ESOC 060801
no no_tm
2006-08-08 06:30:00 32400 1 p3 latched at -68 V. no no_p3
2006-08-08 15:30:00 7200 1 Bias to -40 nA. no bad_bias
2006-08-08 16:57:00 10 3 Bad spacecraft housekeeping packet. no bad_data
2006-08-15 02:31:00 2460 4 WEC off (presumably by DWP voltage monitoring).
1. ESOC 060815
no no_tm
2006-08-21 18:04:07 2 4 Bad packet. no bad_data
2006-08-22 09:15:00 5400 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-08-23 13:09:00 59970 1 Problem with the data (looks digital).
no bad_tm
2006-08-25 05:11:00 7200 1 WEC off by DWP voltage monitoring.
1. ESOC 060825
2. Keith 060825
no no_tm
2006-08-27 12:15:34 2 3 Bad packet. no bad_data
2006-08-27 20:55:00 230400 3 WEC off by DWP voltage monitoring, combined with DWP memory latchup.
1. ESOC 060829
2. ESOC 060830
no no_tm
2006-08-30 12:24:54 43200 3 WEC power cycle to clear DWP memory latchup.
1. ESOC 060915
no no_tm
2006-08-30 15:15:54 44520 2 WEC problem.
1. ESOC 060901
no bad_data
2006-08-31 03:37:00 73380 2 Bias defaults to -40 nA. no bad_bias
2006-09-05 08:14:49 3 1 Bad packet. no bad_data
2006-09-05 15:35:00 6000 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-09-08 17:00:00 46800 2 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-10 17:00:00 81000 2 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-10 19:00:00 68400 3 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-10 21:00:00 63000 4 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-12 16:50:00 8700 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-09-13 00:00:00 79200 3 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-13 02:00:00 79200 2 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-13 02:00:00 79200 4 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-15 09:00:00 79200 3 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-15 10:30:00 79200 4 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-15 13:30:00 477000 1 SC1 payload turned off for 5.5 days with long eclipses, because of bad battery status.
1. ESOC 060718
yes no_tm
2006-09-15 19:30:00 16200 2 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-16 14:12:49 1 3 Bad packet. no bad_data
2006-09-16 14:54:52 33 3 Bad packets. no bad_data
2006-09-16 15:03:47 1 3 Bad packet. no bad_data
2006-09-16 15:07:10 1 3 Bad packet. no bad_data
2006-09-16 15:33:11 1 3 Bad packet. no bad_data
2006-09-16 15:39:01 1 3 Bad packet. no bad_data
2006-09-16 15:54:49 1 3 Bad packet. no bad_data
2006-09-16 16:11:00 1080 3 Bad packets. no bad_data
2006-09-16 16:42:29 1 3 Bad packet. no bad_data
2006-09-17 12:44:00 7560 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-09-18 00:00:00 28800 3 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-18 01:30:00 28800 4 Payload off in long eclipse.
1. ESOC 060718
yes no_tm
2006-09-19 23:06:18 31342 3 Bad EFW data (looks like digital problem).
1. ESOC 060927
no bad_tm
2006-09-27 03:30:00 9000 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-09-30 22:44:45 15 4 Bad HK packet(s). no bad_data
2006-10-03 09:05:00 142320 1 WEC off by DWP voltage monitoring.
1. Per-Arne 061005
2. ESOC 061005
no no_tm
2006-10-05 10:40:40 20 1 Bad HK packet(s). no bad_data
2006-10-08 22:05:00 10800 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-10-12 16:55:58 101069 2 No EFW TM from 16:55:58; all WEC off by DWP voltage monitoring from 20:56. Some details in mail from Keith, apparently with some confusion of times and s/c number.
1. Keith 061013
2. ESOC 061031
no no_tm
2006-10-16 18:29:55 420 2 Maneuver. yes bad_data
2006-10-16 21:00:00 300 1 Maneuver. yes bad_data
2006-10-17 08:30:05 420 3 Maneuver. yes bad_data
2006-10-17 15:29:46 300 4 Maneuver. yes bad_data
2006-10-23 06:10:00 5400 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-10-23 13:32:32 141 3 DWP reset interrupted by scheduled commands caused some confusion and a data gap.
1. Keith 061023
no no_tm
2006-10-23 13:34:53 5167 3 Following the DWP reset confusion described above, EFW operated with bias -40 nA until 15:01 (and then with -140 nA until reset around 061024 17:50, see below). no bad_data
2006-10-23 15:01:00 96648 3 Following the DWP reset confusion described above, EFW operated with bias -40 nA until 15:01, and then with -140 nA until reset around 061024 17:50. no bad_bias
2006-10-27 10:13:38 256800 2 WEC off by DWP voltage monitoring.
1. ESOC 061030
no no_tm
2006-10-28 17:34:00 64740 3 DWP application buffer overflow.
1. ESOC 061102
no TBD
2006-10-29 11:45:00 130500 1 Probe latchup: p3 to -68 V. no no_p3
2006-11-01 10:53:16 4 3 Bad HK packet. no bad_data
2006-11-01 11:08:28 9 3 Bad packet. no bad_data
2006-11-05 12:20:00 104640 2 DWP application buffer overflow.
1. ESOC 061116
no TBD
2006-11-07 03:58:32 2 3 Bad packet. no bad_data
2006-11-11 12:00:00 40 3 Corrupt telemetry. no bad_tm
2006-11-11 12:00:00 90066 3 Probe 2 in Langmuir mode at -40 V for continuous measurement of photosaturation current. Sweep interval 30 minutes.
1. Anders 061020
yes no_p32
2006-11-15 18:20:28 104640 1 S/c problem: SSR controller failure.
1. ESOC 061120
2. AR-99
no TBD
2006-11-16 01:35:00 2460 1 3 V2V34M instead of V23V34M in BM1. yes no_p32
2006-11-18 05:38:00 33120 2 EFW TM stuck, presumably DWP problem.
1. Per-Arne 061125
2. Keith 061127
no bad_tm
2006-11-23 04:39:00 300 1 Maneuver. yes bad_data
2006-11-23 06:35:50 300 2 Maneuver. yes bad_data
2006-11-28 10:37:00 2100 4 WEC power cycle to clear up DWP problem.
1. ESOC 061128
no no_tm
2006-11-29 03:54:05 3 3 Bad packet. no bad_data
2006-12-03 13:48:00 1140 2 DWP application buffer overflow -- may or may not have given EFW problems.
1. ESOC 061206
no TBD
2006-12-03 13:48:00 30300 3 DWP application buffer overflow -- may or may not have given EFW problems.
1. ESOC 061206
no TBD
2006-12-04 13:48:01 30421 3 S/c problem: OBDH anomaly.
1. AR-103
no TBD
2006-12-07 11:16:50 300 1 Maneuver. yes bad_data
2006-12-09 03:09:00 34440 3 Probe latchup. All probes to -68V. no bad_data
2006-12-11 03:30:00 96030 3 Bad burst dump followed by NULL TM mode.
1. Per-Arne 061215
no bad_tm
2006-12-13 10:45:00 3600 4 Spacecraft problem (radiation event). Time is approximate.
1. AR-105
no no_tm
2006-12-13 12:59:00 32460 2 DWP application buffer overflow -- may or may not have given EFW problems.
1. ESOC 061220
no TBD
2006-12-14 00:44:31 5 3 Bad HK. no bad_data
2006-12-15 14:43:58 8 2 Bad HK. no bad_data
2006-12-17 14:51:32 18660 1 WEC switched off by DWP voltage monitoring.
1. ESOC 061217
no no_tm
2006-12-17 23:06:12 8 3 Bad packet. no bad_data
2006-12-17 23:24:11 3 3 Bad packet. no bad_data
2006-12-19 12:05:51 1800 3 WEC power cycle to clear up DWP problem.
1. ESOC 061219
2. Keith 061219
no no_tm
2006-12-20 00:33:00 92220 1 S/c SSR problem. Some data may be available from the real time coverage 18:02 -- 21:38.
1. AR-106
no no_tm
2006-12-22 16:36:24 20160 1 S/c problem: CTU switchover.
1. Keith 061223
2. AR-107
no bad_tm
2006-12-23 10:50:09 95821 2 WEC problem: bad TM until 061224 07:56, when WEC switched off.
1. Per-Arne 061225
no bad_tm
2006-12-26 11:27:40 18000 3 EFW switched off for unknown reason. Switched on again at scheduled reset.
1.
no no_tm
2006-12-29 10:33:00 55620 2 Bias at -40 nA. no bad_bias
2007-01-03 20:37:09 131 2 Corrupted EFW packets. no bad_tm
2007-01-10 17:30:46 6 3 Bad packet. no bad_data
2007-01-11 11:18:00 14160 1 S/c test: switchover to CTU A.
1. ESOC 070111
yes no_tm
2007-01-11 14:34:54 300 2 Maneuver. yes bad_data
2007-01-11 16:34:50 300 4 Maneuver. yes bad_data
2007-01-16 22:54:00 5340 2 WEC switched off by DWP voltage monitoring.
1. ESOC 070117
no no_tm
2007-01-19 03:39:55 5 3 Bad packet. no bad_data
2007-02-02 02:10:00 12000 3 Probes latched at -68 V. no bad_data
2007-02-02 05:26:00 19680 3 WEC switched off by DWP voltage monitoring.
1. ESOC 070202
no no_tm
2007-02-20 06:19:00 10200 4 Strange timing in the EFW data. DWP digital problem? no TBD
2007-02-20 19:02:52 141548 4 Corrupt DWP packets. Data gap until 2007-02-22 10:22. no bad_tm
2007-02-25 05:59:04 113456 4 Data is incomprehensible after DWP watchdog reset. Perhaps EFW was running in a strange mode? Cured by planned reset.
1. ESOC 070226
no bad_data
2007-02-26 16:08:00 80880 1 Bias at -40 nA. no bad_bias
2007-02-28 01:18:34 70320 1 S/c problem: SSR failure.
1. ESOC 070302
2. AR-109
no no_tm
2007-03-01 19:33:08 2 2 Bad packet no bad_data
2007-03-05 12:27:00 3840 3 Payload off in eclipse. yes no_tm
2007-03-06 08:27:23 17576 1 TM buffer overflow.
1. ESOC 070316
no no_tm
2007-03-07 21:31:00 3960 3 Payload off in eclipse. yes no_tm
2007-03-10 06:32:00 4620 1 Payload off in eclipse. yes no_tm
2007-03-10 06:35:00 4440 3 Payload off in eclipse. yes no_tm
2007-03-12 15:37:00 4020 1 Payload off in eclipse. yes no_tm
2007-03-15 00:41:00 3780 1 Payload off in eclipse. yes no_tm
2007-03-17 09:48:00 4620 1 Payload off in eclipse. yes no_tm
2007-03-19 18:56:00 4200 1 Payload off in eclipse. yes no_tm
2007-03-22 04:03:00 4140 1 Payload off in eclipse. yes no_tm
2007-03-22 04:03:00 4020 3 Payload off in eclipse. yes no_tm
2007-03-24 09:53:52 29040 2 EFW switched off, apparently by mistake.
1. Per-Arne 070324
no no_tm
2007-03-24 13:08:00 4860 1 Payload off in eclipse. yes no_tm
2007-03-24 13:33:11 4 4 Bad packet. no bad_data
2007-03-26 22:14:00 3720 3 Payload off in eclipse. yes no_tm
2007-03-26 22:15:00 3960 1 Payload off in eclipse. yes no_tm
2007-03-29 07:23:00 3840 1 Payload off in eclipse. yes no_tm
2007-03-30 07:27:20 21160 3 Probe latchup. p3 to -68V, p2 and 4 to +23 V no bad_data
2007-04-03 19:20:00 44100 3 Probe latchup. p1, p4 go up, p2 goes down. no bad_data
2007-04-04 14:53:20 177000 1 WEC switched off by DWP voltage monitoring at 2007-04-05 15:37:06, but EFW was off since the day before because of a bad reconfiguration following a DWP watchdog reset.
1. ESOC 070406
2. Per-Arne 070412
3. ESOC 070413
no no_tm
2007-04-14 07:47:00 1560 3 S/c problem: corrupted VC4 frames during SSR dump, causing loss of science data for this period.
1. ESOC 070425
no no_tm
2007-04-17 13:15:00 17100 1 Bias at pre-2006 setting (-135 nA). no bad_bias
2007-04-18 09:29:52 300 2 Maneuver. yes bad_data
2007-04-18 10:00:23 300 4 Maneuver. yes bad_data
2007-04-18 10:59:52 300 1 Maneuver. yes bad_data
2007-04-18 11:29:52 480 3 Maneuver. yes bad_data
2007-04-19 22:19:00 46440 1 WEC switched off by voltage monitoring.
1. WecOps 070420
no no_tm
2007-04-23 05:53:29 2 1 Bad packet. no bad_data
2007-04-29 10:16:34 2 3 Bad packet. no bad_data
2007-04-30 21:06:00 24660 2 WEC OFF on SC2 from 120.21.06 to 121.03.58. Cluster Anomaly AR-114 SC2: "CTU Switchover caused by EDAC Double Error".
1. ESOC 070510
2. AR-114
no no_tm
2007-04-30 22:42:10 69260 1 Probe latchup. no bad_data
2007-05-01 09:46:36 3 3 Bad packet. no bad_data
2007-05-04 22:30:00 3600 4 Bad telemetry. Only sporadic packets available. no bad_data
2007-05-06 22:16:00 60420 3 No EFW TLM received + EFW sampling errors received on SC3 from 126.22.16 to 127.15.03.
1. ESOC 070510
no no_tm
2007-05-12 22:19:00 33480 4 WEC switched OFF on SC4 on DOY 132 at 22.19 by the DWP voltage monitoring entry 444. The WEC switch ON was successfully executed from 133.06.31 to 133.07.37.
1. ESOC 070516
no no_tm
2007-05-13 03:23:48 -1 2 Hardware failure: probe 1 dead from this time.
1. Per-Arne 070530
2. Per-Arne 070601
no no_p1
2007-05-13 03:23:59 -1 2 Probe 1 effectively lost from this time.
1. Per-Arne 070525
2. Anders 070530
3. Anders 070604
no no_p1
2007-05-13 22:19:00 32880 4 WEC switched OFF by DWP voltage monitoring entry 444
1. WecOps 070514
no no_tm
2007-05-14 08:12:00 540 1 SC raindance procedure.
1. ESOC 070514
no no_tm
2007-05-18 00:46:00 36660 3 The DWP voltage monitoring entry 445 (raw value "FF") triggered on SC3 on DOY138 at 00.46. The nominal WEC monitoring items configuration were restored on DOY138 at 10.57.
1. ESOC 070524
no no_tm
2007-05-20 12:17:00 30480 3 Application Buffer overflow message received on SC3 from 140.12.17 to 140.20.45.
1. ESOC 070524
no TBD
2007-05-24 22:22:32 180 3 Maneuver yes bad_data
2007-05-25 12:19:35 1682 2 WEC power cycle executed on SC2 on DOY 145 from 12.19.35 to 12.47.37.
1. ESOC 070524
no no_tm
2007-05-28 15:21:00 60 1 3 Strange data at switch to Langmuir mode. yes bad_data
2007-05-28 15:21:00 33486 1 3 Langmuir mode in solar wind, including a 8x450 Hz burst.
1. Anders 070503
yes no_p32
2007-05-28 15:21:00 33486 1 3 Langmuir mode in solar wind, including a 8x450 Hz burst.
1. Anders 070503
yes no_p32
2007-05-29 10:10:00 746 2 P1 diagnostics for hardware failure.
1. Per-Arne 070529
yes no_p1
2007-05-30 06:30:00 744 2 Further P1 diagnostics for hardware failure. Note commanding error for burst test: we used mux position rather than burst list position as argument, resulting in bad burst data. The burst test was rerun on 2007-06-01.
1. Per-Arne 070530
yes no_p1
2007-06-05 13:44:14 75198 3 Only sporadic EFW packets correctly unpacked by TED. EFW-DWP communication problem?
1. Per-Arne 070614
2. Keith 070614
yes bad_tm
2007-06-07 02:41:42 412 2 SC SSR record pointer change.
1. ESOC 070618
no no_tm
2007-06-13 07:51:56 162 2 Further P1 diagnostics for hardware failure. EFW P2 in Langmuir mode, WHI/WBD switching Ey/Ez antennas every few spins, STAFF SA despin off. The test as a whole starts 07:49:36.
1. Anders 070604
2. Jolene 070605
3. Keith 070606
4. ESOC 070611
5. Command sequence (mail attachement)
6. Jolene 070613
7. Anders 070613
8. Jolene 070614
9. Anders 070615
yes no_p1
2007-06-15 18:47:00 51000 1 WEC switched off by DWP voltage monitoring.
1. ESOC 070618
2. ESOC 070621
no no_tm
2007-06-19 07:56:00 540 1 SC raindance procedure.
1. ESOC 070619
no no_tm
2007-06-24 06:23:00 99840 3 Bias at -40 nA. no bad_bias
2007-06-25 10:10:04 79320 2 WEC switched off by DWP voltage monitoring.
1. ESOC 070626
no no_tm
2007-07-04 13:25:24 7722 1 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2007-07-05 12:01:24 39462 1 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2007-07-06 01:19:30 2 1 Bad packet. no bad_data
2007-07-11 05:47:00 3420 3 WEC reconfiguration following DWP reset resulted in default (-40 nA) bias.
1. ESOC 070720
no bad_bias
2007-07-12 06:33:00 4320 1 WEC switched off by DWP voltage monitoring.
1. ESOC 070712
no no_tm
2007-07-21 03:24:07 3450 4 WEC switched off by DWP voltage monitoring.
1. ESOC 070723
no no_tm
2007-07-21 18:00:00 2700 4 Digital corruption. no bad_data
2007-07-21 18:45:00 1235760 2 Bad puck and guard settings (0 V) on C2 for 2007-07-21 18:45 until 2007-08-05 01:01. no TBD
2007-07-25 04:47:00 180 2 Bad data in the LX stream no bad_data
2007-07-25 06:29:52 420 1 Maneuver yes bad_data
2007-07-25 12:59:56 420 3 Maneuver yes bad_data
2007-07-25 13:59:51 300 4 Maneuver yes bad_data
2007-07-29 02:06:47 5 3 Bad HK data no bad_data
2007-07-31 06:47:24 3336 1 WEC switched off by DWP voltage monitoring.
1. ESOC 070731
2. ESOC 070823
no no_tm
2007-08-02 06:28:58 10142 2 WEC switched off by DWP voltage monitoring.
1. ESOC 070802
2. ESOC 070823
no no_tm
2007-08-02 10:29:53 300 1 Maneuver yes bad_data
2007-08-04 07:22:00 63060 4 WEC TM buffer overflow.
1. ESOC 070823
no TBD
2007-08-04 18:20:00 23580 2 WEC TM buffer overflow.
1. ESOC 070823
no TBD
2007-08-04 18:50:00 21780 1 WEC TM buffer overflow.
1. ESOC 070823
no TBD
2007-08-04 21:20:00 12780 3 WEC TM buffer overflow.
1. ESOC 070823
no TBD
2007-08-07 22:35:24 897 2 SC problem: controller failure.
1. ESOC 070808
no no_tm
2007-08-10 12:28:56 745 3 SC raindance procedure.
1. ESOC 070810
no no_tm
2007-08-26 23:35:00 55020 4 Bias at default -40 nA. no bad_bias
2007-08-28 20:06:04 168260 4 Long data gap (until 2007-08-30 18:50:24) caused by CTU switchover.
1. ESOC 070831
no no_tm
2007-08-31 17:42:28 50312 2 Mixed-up quantities in hx telemetry. no bad_data
2007-09-01 00:00:00 27720 2 Bias at -40 nA. no bad_bias
2007-09-03 23:57:40 5 4 Bad HK packet. no bad_data
2007-09-16 19:10:22 11688 3 Probe latchup. no bad_data
2007-09-26 19:10:00 62400 2 Mixed-up quantities in HX telemetry. no no_p34
2007-09-26 19:10:00 62400 2 Bias at -40 nA. no bad_bias
2007-09-26 19:10:00 5700 2 Strange p3 data. Likely related to bad bias (see above). no no_p3
2007-09-29 10:23:00 25576 4 ESOC data loss.
1. ESOC 071203
no no_tm
2007-09-30 23:36:00 2302 4 ESOC data loss.
1. ESOC 071203
no no_tm
2007-10-12 04:30:00 36000 4 Choppy telemetry. no TBD
2007-10-13 07:21:48 72741 1 Digital hickup in EFW: quantities badly mixed up.
1. Anders 081202
no bad_tm
2007-10-15 14:52:02 78478 4 Probe latchup (all probes -68) no bad_data
2007-10-21 06:40:00 420 2 DWP command response failure.
1. Keith 071105
no bad_tm
2007-10-22 11:13:00 420 2 DWP command response failure.
1. Keith 071105
no bad_tm
2007-10-22 13:35:00 420 2 DWP command response failure.
1. Keith 071105
no bad_tm
2007-10-24 19:29:50 420 3 Maneuver. yes bad_data
2007-10-24 22:59:53 420 1 Maneuver. yes bad_data
2007-10-28 11:14:15 5 1 Bad HK packet. no bad_data
2007-10-30 18:49:00 6060 4 Long pulsed radial delta-v maneuver.
1. ESOC MOR
no bad_data
2007-11-02 02:05:24 8142 1 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2007-11-02 06:00:00 5460 2 Long pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-03 00:18:29 3 3 Bad packet. no bad_data
2007-11-04 12:50:00 4140 1 Long pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-04 14:21:10 5940 3 Long pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-05 15:22:38 1202 3 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-08 03:04:10 1685 4 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-10 03:29:35 78325 2 Bias to default -40 nA, followed by a period of -135 nA (until 22:00). no bad_bias
2007-11-10 08:55:09 60 3 Single pulse for radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-10 09:58:48 1298 1 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-10 10:57:17 1537 2 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-16 10:43:04 5700 4 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-16 12:20:00 4800 1 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-16 13:40:00 4000 2 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-18 17:36:56 5464 3 Pulsed radial delta-v maneuver.
1. ESOC MOR
yes bad_data
2007-11-22 05:23:06 260 3 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-22 05:46:14 15 1 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-22 06:30:04 8 4 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-22 06:51:56 20 2 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-23 01:10:00 7200 4 Serious telemetry errors. Likely WEC glitch. no bad_data
2007-11-23 14:50:00 45 1 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-23 14:51:00 120 4 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-23 16:07:38 325 2 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-23 18:12:00 155 3 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-29 09:03:59 5 1 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-29 09:43:03 5 4 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-29 10:03:02 5 2 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-29 12:00:00 60 3 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-11-30 15:27:31 45 3 Pulsed radial delta-v trim plus spin-up.
1. ESOC MOR
yes bad_data
2007-12-04 03:53:43 270 3 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-12-06 12:24:02 60 1 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-12-06 13:01:22 5 4 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-12-06 13:18:34 5 2 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-12-08 22:05:45 15 3 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-12-13 16:14:03 5 3 Pulsed radial delta-v trim.
1. ESOC MOR
yes bad_data
2007-12-19 21:00:00 95400 4 Bias at -40 nA. no bad_bias
2007-12-25 12:45:00 37440 2 Choppy telemetry. WEC TM overflow due to SEU.
1. Keith 080108
no bad_data
2008-01-04 09:07:00 30660 1 Corrupt phase data. no bad_data
2008-01-14 14:42:37 23939 2 WEC Switched off
1. WEC Ops 080115
no no_tm
2008-01-17 03:18:47 50268 1 WEC Switched off
1. WEC Ops 080117
no no_tm
2008-01-22 06:11:02 120238 4 Probe latchup.
1. Per-Arne 080123
no bad_data
2008-01-24 06:11:22 28538 1 WEC Switched off
1. WEC Ops 080124
no no_tm
2008-01-28 12:00:00 390 2 Maneuver
1. ESOC MOR
yes bad_data
2008-01-28 13:30:00 390 3 Maneuver
1. ESOC MOR
yes bad_data
2008-01-28 15:00:00 390 4 Maneuver
1. ESOC MOR
yes bad_data
2008-01-28 16:30:00 390 1 Maneuver
1. ESOC MOR
yes bad_data
2008-01-31 20:48:00 52020 3 Bias at -40 nA. no bad_bias
2008-02-04 15:00:00 5 1 2 3 4 Trim maneuvers on all 4 sats.
1. ESOC MOR
yes bad_data
2008-02-08 22:16:30 44490 2 Data gap due to CTU switchover (see next entry). no no_tm
2008-02-09 15:04:50 166510 2 CTU switchover resulted in data loss.
1. ESA 080212
no no_tm
2008-02-17 01:46:00 120480 2 Data gap related to WEC latchup at 2008-02-12T05:43:55Z.
1. Keith 082018
no no_tm
2008-02-25 07:50:00 203100 3 Multiple isolated telemetry glitches (spikes in E data). Also, sending p12 rather than p34 until 02-26 14:29. no TBD
2008-02-27 22:31:33 1 2 Bad packet. no bad_data
2008-02-27 22:36:35 1 2 Bad packet. no bad_data
2008-03-11 05:50:00 59280 1 Corrupt telemetry. no bad_tm
2008-03-14 13:28:00 85320 2 SSR failure documented in CLU_AR-130.
1. MOR 397
no no_tm
2008-03-16 12:32:31 2 4 Strange spike in the p3/p4 data not seen in p1/p2. no bad_data
2008-03-26 23:51:38 1 1 Possible battery gas release.
1. Per-Arne 0800411
no bad_data
2008-04-18 00:53:00 79320 4 WEC monitoring 445 triggered switchover to default (-40 nA) bias.
1. Keith 080421
no bad_bias
2008-04-19 08:34:00 177960 1 WEC latchup.
1. Keith 080421
1. ESOC 080421
no no_tm
2008-04-22 01:37:00 27480 4 Repeating VC2 led to corruption in SSR.
1. Juergen 080425
no no_tm
2008-04-23 00:10:44 1 2 Spike in p32. no bad_data
2008-04-23 02:51:00 51540 1 EFW probe latchup.
1. Per-Arne 0800428
no bad_data
2008-04-25 04:58:00 0 3 Error in ADCE caused switchover to backup ADCE-B.
1. Juergen 080425
no TBD
2008-04-25 09:00:00 960 3 Maneuver.
1. ESOC MOR
yes bad_data
2008-04-25 10:30:00 240 4 Maneuver. Time in MOR is incorrectly given as 20:30 rather than 10:30.
1. ESOC MOR
yes bad_data
2008-04-25 12:00:00 240 1 Maneuver.
1. ESOC MOR
yes bad_data
2008-04-25 13:00:00 240 2 Maneuver.
1. ESOC MOR
yes bad_data
2008-04-25 15:14:00 13740 2 TM Buffer Overflow
1. Colette 080426
no no_tm
2008-04-25 17:48:00 4500 1 3 TM Buffer Overflow
1. Colette 080426
no no_tm
2008-04-28 08:00:00 360 3 Maneuver (C3 tilt).
1. ESOC MOR
yes bad_data
2008-04-30 21:00:00 20 3 Maneuver (C3 tilt).
1. ESOC MOR
yes bad_data
2008-05-07 09:35:55 35400 4 Repeating VC2 VCFC caused data loss.
1. ESOC MOR
1. Juergen 080514
no no_tm
2008-05-07 12:43:25 36000 3 DWP upset
1. Keith 080605
no no_tm
2008-05-07 23:30:00 5 3 Maneuver (C3 tilt).
1. ESOC MOR
yes bad_data
2008-05-10 18:16:38 136665 1 SSR anomaly
1. Juergen 080514
no no_tm
2008-05-12 12:14:03 120593 1 SSR anomaly
1. Juergen 080514
no no_tm
2008-05-14 08:00:00 360 3 Maneuver (C3 tilt).
1. ESOC MOR
yes bad_data
2008-05-17 02:54:00 18060 1 SSR anomaly
1. Juergen 080520
no no_tm
2008-05-17 11:24:00 31680 1 SSR anomaly
1. Juergen 080520
no no_tm
2008-05-20 00:20:00 16800 3 Data gap affecting all instruments. May be explained in MOR 407 (missing). no no_tm
2008-05-29 07:22:00 8400 1 EFW loss of TM because of spurious switch to NULL mode.
1. Per-Arne 080529
no bad_data
2008-05-30 08:00:00 1440 3 Maneuver (C3 tilt).
1. ESOC MOR
yes bad_data
2008-06-13 06:23:30 37230 3 Probe latchup no bad_data
2008-06-18 13:31:46 5 3 Maneuver.
1. ESOC MOR
yes bad_data
2008-06-20 19:55:04 5 4 Bad HK packet. no bad_data
2008-06-24 15:19:00 96360 1 EFW spurious switch-off, later followed by WEC switch-off.
1. Per-Arne 080625
no no_tm
2008-06-24 21:33:00 57420 4 Default -40 nA bias no bad_bias
2008-06-27 08:26:43 2 3 Bad HK packet. no bad_data
2008-07-01 18:29:00 78660 1 DWP reset by onboard monitoring. WEC in emergency mode.
1. Keith 080702
no bad_bias
2008-07-03 09:52:29 53940 3 EFW switched off by DWP upset.
1. Keith 080704
no no_tm
2008-07-04 00:18:23 4 1 Bad packet no bad_data
2008-07-04 00:38:19 4 1 Bad packet no bad_data
2008-07-05 20:16:00 1500 4 Bad phase from HK no bad_data
2008-07-12 04:10:28 90000 1 WEC off by voltage monitoring.
1. ESOC 080713
no no_tm
2008-07-13 23:30:08 5 1 Bad HK packet no bad_data
2008-07-18 14:27:08 26 3 Bad packet no bad_data
2008-07-18 14:29:00 16 3 Bad phase from HK no bad_data
2008-07-28 22:01:35 90 2 Manoeuvre (16 pulses) yes bad_data
2008-07-28 22:42:31 90 1 Manoeuvre (18 pulses) yes bad_data
2008-07-28 23:39:46 30 3 Manoeuvre (7 pulses) yes bad_data
2008-07-30 12:30:00 240 2 Manoeuvre (56 pulses) yes bad_data
2008-07-30 13:30:00 240 1 Manoeuvre (55 pulses) yes bad_data
2008-07-30 14:30:00 280 3 Manoeuvre (63 pulses) yes bad_data
2008-07-30 15:30:00 260 4 Manoeuvre (58 pulses) yes bad_data
2008-07-31 06:26:25 32220 3 EFW switched off by DWP upset.
1. Keith 080812
no no_tm
2008-07-31 15:15:28 91809 1 WEC off by voltage monitoring.
1. Juergen 080801
no no_tm
2008-08-02 16:58:27 5 1 Manoeuvre (1 pulse) yes bad_data
2008-08-02 17:58:10 5 3 Manoeuvre (1 pulse) yes bad_data
2008-08-02 18:45:57 5 4 Manoeuvre (1 pulse) yes bad_data
2008-08-06 00:16:00 49440 3 Default bias (-40 nA). no bad_bias
2008-08-08 17:04:24 21222 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-08-14 16:40:24 119862 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-08-15 10:32:24 55542 1 2 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-08-20 13:31:24 23382 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-08-23 07:06:00 280380 1 EFW off for unknown reason; all WEC swithched off by onboard monitoring due to high EFW inrush current when a scheduled reset tried to start EFW with all WEC already on at 2008-08-25 13:56:08.
1. Per-Arne 080825
2. Per-Arne 080826
no no_tm
2008-08-23 14:26:00 172800 3 Default bias (-40 nA) for 2 days. no bad_bias
2008-08-26 19:40:24 98562 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-09-02 01:35:53 2 3 Bad packet no bad_data
2008-09-04 08:50:40 5 1 Bad HK packet no bad_data
2008-09-08 23:43:00 103140 3 Default bias (-40 nA). no bad_bias
2008-09-16 11:18:57 3 3 Bad packet no bad_data
2008-09-21 13:35:18 3 3 Bad packet no bad_data
2008-09-22 03:15:00 15120 4 Burst (BM1) data replaced by NM data for all of WEC because of DWP problem.
1. Keith 081010
no TBD
2008-09-22 12:28:27 172893 3 WEC switched off after inrush current triggered monitor.
1. Keith 081022
1. Colette 081023
1. Juergen 081023
no no_tm
2008-09-25 18:40:24 52962 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-09-29 12:05:48 61998 1 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-10-04 05:15:24 24642 1 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-10-04 11:58:00 8400 1 WEC switched off by voltage monitoring 444
1. Colette 081006
no no_tm
2008-10-08 21:18:30 90 2 Large swing in Vsc, presumably due to EDI.
1. pal/aie 100426
no TBD
2008-10-12 11:28:00 67410 3 numerous short data gaps in EFW
1. Per-Arne 081028
no bad_data
2008-10-13 06:51:24 102762 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-10-20 18:34:00 208440 1 2 3 4 Sweep interval 30 minutes on all s/c for one orbit.
yes TBD
2008-10-22 04:46:00 28000 3 EFW stops working, but continues to present data.
1. Per-Arne 100426
no bad_data
2008-10-24 02:42:00 417600 1 EFW digital problem: instrument stopped working, recovering at next reset. (dt is approximate)
1. Per-Arne 081027
2. ESOC 081028
no no_tm
2008-10-27 01:02:35 309000 3 DWP watchdog reset caused WEC off.
1. ESOC 081121
no no_tm
2008-11-01 20:50:06 2 3 Bad packet no bad_data
2008-11-04 14:40:24 20382 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-11-10 16:09:12 50 2 Manoeuvre (10 pulses) yes bad_data
2008-11-10 16:21:38 60 1 Manoeuvre (12 pulses) yes bad_data
2008-11-10 16:40:08 180 3 Manoeuvre (39 pulses) yes bad_data
2008-11-11 08:06:24 8622 1 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-11-11 13:00:00 300 1 Manoeuvre (67 pulses, attitude) yes bad_data
2008-11-11 14:00:00 300 2 Manoeuvre (68 pulses, attitude) yes bad_data
2008-11-11 15:00:00 410 3 Manoeuvre (90 pulses, attitude) yes bad_data
2008-11-11 16:30:40 360 4 Manoeuvre(71 pulses, attitude) yes bad_data
2008-11-15 11:37:36 40 3 Manoeuvre (8 pulses) yes bad_data
2008-11-15 13:02:14 928 4 Data loss due to CTU switchover (see below)
1. Juergen 081124
no no_tm
2008-11-15 15:42:00 84180 4 Data loss due to CTU switchover (see below)
1. Juergen 081124
no no_tm
2008-11-16 18:15:00 32880 4 Data loss due to CTU switchover (see below)
1. Juergen 081124
no no_tm
2008-11-17 08:49:55 21000 4 Data loss due to CTU switchover caused by corrupted command upload.
1. Juergen 081117
1. Juergen 081118
no no_tm
2008-11-27 09:05:22 90 1 Manoeuvre (18 pulses) yes bad_data
2008-11-27 09:15:16 5 4 Manoeuvre (1 pulse) yes bad_data
2008-11-27 15:30:01 10 3 Manoeuvre (2 pulses) yes bad_data
2008-11-27 16:20:02 140 2 Manoeuvre (28 pulses) yes bad_data
2008-11-28 03:15:33 3214 2 S/c problem: SSR pointer corruption.
1. AR-128
no no_tm
2008-12-02 02:47:07 5 1 Manoeuvre (1 pulse) yes bad_data
2008-12-02 02:51:31 5 2 Manoeuvre (1 pulse) yes bad_data
2008-12-02 05:33:13 10 3 Manoeuvre (2 pulses) yes bad_data
2008-12-03 07:15:10 50700 1 WEC switched off after a latchup in the DWP interrupted normal bootup.
1. Keith 081204
no no_tm
2008-12-06 20:58:44 5 1 Manoeuvre (1 pulse) yes bad_data
2008-12-06 21:03:40 5 2 Manoeuvre (1 pulse) yes bad_data
2008-12-06 21:05:00 5 4 Manoeuvre (1 pulse) yes bad_data
2008-12-06 23:45:46 10 3 Manoeuvre (2 pulses) yes bad_data
2008-12-09 07:25:24 9582 2 3 118 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2008-12-09 18:16:00 59400 4 Bias at default -40 nA no bad_bias
2008-12-11 17:55:05 5 3 Manoeuvre (1 pulse) yes bad_data
2008-12-20 04:47:23 98400 4 WEC switched off after memory latchup.
1. Keith 081223
no no_tm
2008-12-22 22:27:52 79320 1 WEC switched off after DWP watchdog reset
1. Keith 081224
no no_tm
2008-12-26 13:44:00 46680 4 Bias at default -40 nA no bad_bias
2008-12-28 03:35:29 50400 1 WEC switched off due to monitoring entries 444/445/446. Powered up in emergency mode (see below).
1. Keith 081229
no no_tm
2008-12-28 17:30:00 17640 1 WEC powered up in emergency mode, with -40 nA bias settings.
1. Keith 081229
no bad_bias
2009-01-06 16:28:34 127200 4 WEC switched off due to monitoring entry 444.
1. ESOC 090108
1. Keith 090108
no no_tm
2009-01-13 15:33:15 23326 3 S/c problem: data lost due to SSR anomaly.
1. ESOC 090115
no no_tm
2009-01-14 01:07:49 23519 3 S/c problem: data lost due to SSR anomaly.
1. ESOC 090115
2. AR-139
no no_tm
2009-01-18 23:06:11 50306 1 WEC switch-off due to DWP memory latchup.
1. Keith 090120
no no_tm
2009-01-22 14:06:00 1200 2 WEC power-cycled to clear latchup. Data may potentially be affected from 2009-01-21 23:44, although no obvious problems.
1. Keith 090122
no no_tm
2009-01-29 13:15:10 2 3 Bad packet. no bad_data
2009-02-04 13:44:00 114300 4 DWP reset by onboard monitor, emergency macro failed.
1. Keith 090205
no no_tm
2009-02-05 11:30:00 360 3 Maneuver (84 pulses)
1. MOR 444
yes bad_data
2009-02-05 12:30:00 300 4 Maneuver (69 pulses)
1. MOR 444
yes bad_data
2009-02-13 02:20:00 52680 1 Bias at default -40 nA. no bad_bias
2009-02-13 09:45:39 629 3 ESOC rain dance procedure.
1. ESOC 090213
yes no_tm
2009-02-15 01:34:08 132420 2 WEC switched off due to onboard monitoring
1. ESOC 090216
no no_tm
2009-02-15 08:31:37 4 1 Bad HK packet no bad_data
2009-02-18 14:23:40 20880 1 EFW telemetry garbled. Affects both HX and LX. Cleared by switch to burst mode at 20:11.
1. Keith 090219
no bad_data
2009-02-18 17:06:11 3900 3 Data loss due to s/c SSR anomaly.
1. ESOC 090224
no no_tm
2009-02-25 16:25:42 32718 3 Probe latchup. no bad_data
2009-03-01 06:23:01 600 4 S/c battery leak.
1. Per-Arne 090303
no bad_data
2009-03-02 17:30:00 86280 4 Default -40 nA bias. no bad_bias
2009-03-08 01:15:49 45092 3 S/c problem: SSR controller failure.
1. AR-141
no no_tm
2009-03-12 16:42:00 68400 1 EFW switched off by DWP for unknown reason; all WEC switched off at 2009-03-13T05:52:00Z.
1. Per-Arne 090313
2. ESOC 090313
no no_tm
2009-03-13 16:11:00 25565 1 EFW data loss presumably due to CTU switchover (see next entry).
1. Per-Arne 090326
no no_tm
2009-03-13 23:17:05 22140 1 S/c problem: CTU/RTU switch-over.
1. ESOC 090314
2. AR-142
no no_tm
2009-03-15 12:29:31 39000 4 WEC turned off by onboard monitoring.
1. Keith 090316
no no_tm
2009-03-18 05:41:02 3 3 bad HX packet no bad_data
2009-03-19 01:01:45 48120 2 DWP commanding error caused first slightly corrupt HX data until 07:55, then highly corrupt data (including bad LX). Interval until 07:55 may be recoverable.
1. Keith 090319
no bad_data
2009-03-21 00:00:24 25782 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-03-31 12:00:00 5580 4 Choppy HX telemetry for this burst. no bad_data
2009-03-31 14:13:00 59220 3 Default -40 nA bias. no bad_bias
2009-04-03 01:38:41 2 1 Bad packet. no bad_data
2009-04-06 08:56:33 57000 3 TM overflow following DWP watchdog reset.
1. Keith 090420
no bad_tm
2009-04-08 04:43:00 19020 3 EFW in TM mode NULL for unknown reason.
1. Per-Arne 090408
no no_tm
2009-04-17 18:01:07 18720 2 Single event upset corrupted the DWP application code for Whisper, caused serious DWP problems.
1. Keith 090430
no bad_data
2009-04-19 07:28:59 1592126 1 P4 failed on C1. It mysteriously came back to life on 2009-05-07, before failing for good on 2009-10-14.
1. Per-Arne 090504
no no_p4
2009-04-22 01:44:00 53640 3 Default -40 nA bias. no bad_bias
2009-04-24 10:53:19 11598 1 EFW powered off following DWP watchdog reset.
1. Keith 090424a
no no_tm
2009-04-29 21:40:27 2 3 Bad packet. no bad_data
2009-04-30 19:56:37 61980 4 Data loss due to CTU switchover.
1. Juergen 090501
no no_tm
2009-05-02 20:40:00 12288 1 Data lost when spacecraft went into eclipse
1. MOR 456
no no_tm
2009-05-06 12:00:00 1260 1 Tests for the broken probe 4.
1. Per-Arne 090507
yes bad_data
2009-05-07 17:44:22 14438 1 Probe 4 returns from the dead for 4 hours and then fails again.
1. Per-Arne 090511
no TBD
2009-05-08 07:30:00 -1 1 Probe 4 functional again. Works until 2009-10-14.
1. Per-Arne 090511
no TBD
2009-05-09 08:15:06 114300 1 P4 switched into density mode (V=0) when it was thought to be dead, then switched back to E-field mode the next day when it was discovered to be working.
1. ESOC 090508
yes no_p4
2009-05-09 17:01:24 58422 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-05-15 11:00:00 270 1 Maneuver (59 pulses)
1. MOR 465
yes bad_data
2009-05-15 12:00:00 280 2 Maneuver (60 pulses)
1. MOR 465
yes bad_data
2009-05-15 13:00:00 370 3 Maneuver (83 pulses)
1. MOR 465
yes bad_data
2009-05-15 14:00:00 290 4 Maneuver (63 pulses)
1. MOR 465
yes bad_data
2009-05-19 03:40:00 10200 2 Corrupt data no bad_data
2009-06-02 14:09:00 75360 3 Data loss due to DWP latchup.
1. Colette 090603
no no_tm
2009-06-02 18:32:00 144000 1 DWP emergency macro caused data loss and some fill data
1. Keith 090605
no bad_data
2009-06-10 17:03:55 66345 3 WEC operations "very disturbed" due to DWP latchup. EFW does not seems affected.
1. Keith 090613
no TBD
2009-06-11 11:29:40 293700 3 WEC switched off due to DWP latchup. Combined with above problem, no C3 data for 4 days.
1. Keith 090613
2. Keith 090615
no no_tm
2009-06-13 12:05:00 857460 1 DWP latchup: EFW appears unaffected.
1. Keith 090623
no TBD
2009-06-19 18:47:00 327180 4 DWP latchup: EFW appears unaffected, except for the interval below.
1. Keith 090623
no TBD
2009-06-20 02:09:00 1234 4 p3 develops a strange offset, and then abruptly returns to normal. p34 data consequently weird. no bad_data
2009-06-24 19:19:30 54000 3 WEC switched off by onboard monitor 445.
1. Keith 090625
2. Colette 090702
no no_tm
2009-06-29 11:22:50 102780 3 EFW telemetry stuck at fixed values, though analog line to other WEC instruments appear good.
1. Per-Arne 090709
no bad_data
2009-07-09 20:03:00 28080 3 Patchy EFW data. Some EFW packets getting incorrectly tagged as STAFF. Incorrectly-tagged data may be recoverable if it's truly essential.
1. Nicole 090811
2. Keith 090825
no bad_data
2009-07-10 12:50:00 5100 4 WEC switched off, probably onboard monitoring.
1. ESOC 090710
no no_tm
2009-07-18 07:46:05 660 2 Corrupt telemetry no bad_data
2009-07-20 00:20:00 18420 1 WEC switched off due to onboard monitoring 01BD, 01BE, 01BC.
1. Jefferys 090720
no no_tm
2009-07-24 09:43:04 3 3 Corrupt telemetry no bad_data
2009-07-28 13:35:00 42900 2 WEC in strange state; ISDAT has a hard time unpacking the data, with gaps that come and go. WEC turned off at 23:33.
1. ESOC 090729
no TBD
2009-08-03 17:55:00 165600 4 Data loss due to s/c problem (CTU switchover).
1. ESOC 090805
2. AR-151
no no_tm
2009-08-04 11:35:00 16200 2 Only 25 Hz data available in BM1 due to DWP problem.
1. Keith 090825
no TBD
2009-08-08 17:31:43 36840 1 WEC switched of by onboard monitoring.
1. ESOC 090809
no no_tm
2009-08-11 03:35:49 213120 3 WEC switched of by onboard monitoring.
1. Keith 090812
2. Keith 090813
3. ESOC 090813
4. Keith 090813
no no_tm
2009-08-28 06:00:00 240 4 Maneuver (52 pulses)
1. MOR 473
yes bad_data
2009-08-28 06:30:00 300 3 Maneuver (70 pulses)
1. MOR 473
yes bad_data
2009-08-28 07:00:00 240 2 Maneuver (51 pulses)
1. MOR 473
yes bad_data
2009-08-28 08:00:00 230 1 Maneuver (49 pulses)
1. MOR 473
yes bad_data
2009-09-01 10:09:24 9342 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-02 17:50:40 121920 1 WEC switched off by watchdog.
1. Colette 090904
2. Keith 090904
no no_tm
2009-09-03 20:49:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-06 04:55:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-08 06:37:24 34242 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-09 01:19:38 198000 3 Long data loss due to WEC watchdog reset caused by WBD command malfunction.
1. Keith 090909
2. Colette 090909
3. Keith 090911
no no_tm
2009-09-13 08:58:24 6342 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-13 14:45:24 99462 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-15 16:52:24 9282 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-18 23:02:00 16080 3 Data lost due to s/c problem (SSR controller failure).
1. AR-152
no no_tm
2009-09-20 11:50:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-20 13:48:48 38400 1 WEC off by monitoring entries 445/446/444.
1. Keith 090921
no no_tm
2009-09-20 22:55:24 5976 1 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-20 22:55:24 82242 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-22 13:00:00 14400 2 WEC switched off for s/c tests.
1. ESOC 090910
yes no_tm
2009-09-22 20:43:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-09-30 10:33:00 44820 1 WEC off by monitoring entries 445/446/444.
1. Jefferys 091001
no no_tm
2009-10-01 09:05:24 3942 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-02 06:26:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-03 08:17:24 39762 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-04 14:58:30 111540 4 WEC operation "very disturbed". EFW data has big spikes and glitches.
1. Colette 091006
no bad_data
2009-10-04 15:09:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-06 02:54:24 3942 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-08 05:00:00 30600 3 Glitchy HX telemetry. no bad_data
2009-10-08 05:52:00 27660 1 EFW in emergency mode (wrong filters and bias) after WEC monitoring 445 triggered.
1. Colette 091015
no bad_data
2009-10-10 20:52:00 2700 1 2 3 Bias mode toggling in support of PEACE photoelectron investigations.
1. Anders 090923
2. Report 091203
yes bad_bias
2009-10-12 08:28:30 83490 1 Corrupt HX and LX telemetry. no bad_data
2009-10-13 01:02:00 2700 1 2 3 Bias mode toggling in support of PEACE photoelectron investigations.
1. Anders 090923
2. Report 091203
yes bad_bias
2009-10-14 04:24:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-14 07:00:00 -1 1 P4 fails again, never to return.
1. Anomaly report CLU_PL-6
no no_p4
2009-10-17 16:41:00 2700 1 2 3 Bias mode toggling in support of PEACE photoelectron investigations.
1. Anders 090923
2. Report 091203
yes bad_bias
2009-10-19 11:28:00 42000 2 Corrupt telemetry (HX and LX) no bad_data
2009-10-19 23:51:00 2700 1 2 3 Bias mode toggling in support of PEACE photoelectron investigations.
1. Anders 090923
2. Report 091203
yes bad_bias
2009-10-21 07:44:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-10-21 19:35:00 21420 1 Spacecraft in nominal survival mode after power-down eclipse.
1. Juergen 091023
no no_tm
2009-10-28 04:01:24 36582 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-11-02 05:10:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2009-11-06 16:30:00 17400 4 Stuck DAC making weird higher-freq signals. Spinfits OK. no info
2009-11-11 07:31:03 8 1 Strange phase from isdat. no bad_data
2009-11-12 13:45:26 33060 3 SSR controller failure
1. Juergen 091117
no no_tm
2009-11-19 15:00:00 270 3 Attitude slew manoeuvre (58 pulses)
1. MOR 485
yes bad_data
2009-11-19 15:45:00 200 1 Attitude slew manoeuvre (42 pulses)
1. MOR 485
yes bad_data
2009-11-19 16:30:00 200 4 Attitude slew manoeuvre (45 pulses)
1. MOR 485
yes bad_data
2009-11-19 17:00:00 200 2 Attitude slew manoeuvre (45 pulses)
1. MOR 485
yes bad_data
2009-11-24 17:41:00 29880 4 Corrupt data after DWP watchdog reset.
1. Per-Arne 091223
2. Keith 091223
no bad_data
2009-11-24 17:43:00 29880 4 Bad timing solved by WEC restart. no bad_data
2009-11-30 13:15:30 80 4 DAC stuck after sweep. no bad_data
2009-12-01 12:00:00 31500 4 WEC in a strange state. ISDAT has a tough time unpacking the HK, although no obvious problems in EFW data. Few details.
1. Colette
2. Keith 091202
no TBD
2009-12-04 05:20:00 429000 4 Probes latched up for 5 days.
1. Per-Arne 091208
no bad_data
2009-12-05 09:07:00 175200 3 Watchdog reset triggered monitoring entry 436 and shut off WEC. 48 hours data lost.
1. Keith 091205
no no_tm
2009-12-08 07:13:10 50400 1 DWP watchdog reset switched off EFW.
1. Keith 091209
no no_tm
2009-12-09 07:58:56 11 1 Bad phase data. no bad_data
2009-12-09 21:48:00 11400 1 Bad timing in WEC packets. no bad_data
2009-12-25 17:33:00 8280 2 WEC problem, cleared by scheduled reset.
1. Per-Arne 100106
no bad_data
2009-12-28 03:05:00 92820 4 WEC switched off by monitoring entry 444.
1. Keith 091231
no no_tm
2010-01-03 18:38:44 20 1 TM Buffer Overflow.
1. Collette 100121
no bad_data
2010-01-06 01:00:58 12 2 TM buffer overflow
1. Collette 100125
no bad_data
2010-01-06 03:35:00 26100 3 SSR sropped recording. Anomaly report.
1. Sylvia 100113
no no_tm
2010-01-06 19:37:50 20 1 TM buffer overflow
1. Collette 100125
no bad_data
2010-01-09 11:07:52 -1 1 2 3 4 Sliding window set to LX (instead of spinfits) from this time.
1. Keith 091217
yes TBD
2010-01-10 05:34:13 43560 2 DWP watchdog reset.
1. Keith 100111
no no_tm
2010-01-10 13:31:00 169800 3 DWP watchdog reset. 48 hours of data lost.
1. Keith 100112
no no_tm
2010-01-12 19:44:34 10 1 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-13 15:00:00 300 2 Manoeuvre. (50 pulses)
1. ESOC MOR
yes bad_data
2010-01-13 15:45:00 300 3 Manoeuvre. (64 pulses)
1. ESOC MOR
yes bad_data
2010-01-13 16:30:00 300 4 Manoeuvre. (50 pulses)
1. ESOC MOR
yes bad_data
2010-01-13 17:15:00 300 1 Manoeuvre. (46 pulses)
1. ESOC MOR
yes bad_data
2010-01-14 14:27:32 10 3 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-14 20:42:17 8 1 Strange phase from isdat. no bad_data
2010-01-15 01:01:02 10 1 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-15 02:02:12 15 2 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-15 02:03:00 120 3 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-15 02:30:00 190800 3 Probes 3 and 4 in a strange state. p2 data may or may not be useful. Commanding error? no bad_data
2010-01-17 01:06:24 12342 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-01-17 16:21:05 10 4 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-17 20:10:00 55200 4 Default -40 nA bias. no bad_bias
2010-01-19 19:48:04 15 3 TM Buffer overflow
1. Collette 100129
no bad_data
2010-01-21 07:35:24 34302 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-01-21 20:48:00 84300 1 WEC off by monitoring 444.
1. ESOC 100122
no bad_data
2010-01-25 16:50:40 70 4 DAC error. no bad_data
2010-01-26 00:21:00 36600 2 SSR anomaly
1. Silvia 100127
no no_tm
2010-01-28 17:01:00 26880 2 Commanding mistake
1. Juergen, Per-Arne 100219
no no_tm
2010-01-30 10:03:24 29322 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-02-01 08:33:00 1500 1 Data loss for all payload.
1. AR-2624
no no_tm
2010-02-01 10:10:00 17940 1 Data loss for all payload.
1. AR-2624
no no_tm
2010-02-01 17:47:00 51720 1 Data loss for all payload.
1. AR-2624
no no_tm
2010-02-04 07:55:00 429900 1 Undervoltage in CTU caused serious payload problems. Roughly 5 days of data lost. Anomaly: CLU-AR-162 SC1 does not accept block commands
1. Juergen 100209
no no_tm
2010-02-06 05:00:24 29082 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-02-12 00:28:32 6 1 Bad phase data no bad_data
2010-02-12 08:50:00 10200 4 SSR anomaly
1. Juergen 100215
no no_tm
2010-02-15 14:15:00 13500 2 3 Instruments operated through eclipse at high bias. no bad_bias
2010-02-24 14:59:55 25 3 Manoeuvre with instruments on.
1. Juergen 100226
no bad_data
2010-03-01 05:53:00 10800 1 Bad TM buffer overflows.
1. Colette 100317
no bad_data
2010-03-02 13:40:46 72560 2 DWP reset caused bad data.
1. Keith 100303
no bad_data
2010-03-04 18:00:00 81240 2 Bias at defualt -40 nA. no info
2010-03-26 10:23:00 45600 4 WEC off after monitoring
1. ESOC 100330
no no_tm
2010-03-26 18:13:00 51000 2 WEC monitor 445. EFW data is corrupt (10 Hz filters, full of spikes).
1. Colette 100414
no bad_data
2010-03-27 05:25:00 9900 1 Unusual burst setup: V43M and SCY sampled at 18 kHz. Caused use of SPLIT sampling and attendant problems with V23 data.
1. SPLIT mode problem
no no_p32
2010-03-30 04:32:18 2 1 Bad packet. no bad_data
2010-04-06 04:40:24 8982 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-07 17:40:24 41382 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-10 16:05:24 17922 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-12 22:40:24 16902 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-13 07:57:45 205200 2 EFW turn-off by WEC following probable DWP SEU.
1. Keith 100414
2. ESOC 100415
3. Keith 100417
no no_tm
2010-04-15 01:40:24 28722 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-16 14:21:00 3000 1 WEC error.
1. Colette 100416
no no_tm
2010-04-17 04:50:24 39822 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-19 16:40:24 19782 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-20 21:46:00 27240 1 anomaly AR-135: repeating VC2 virtual channelframe counter on C1
1. Juergen 100423
no no_tm
2010-04-21 23:15:24 18702 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-23 23:36:24 39822 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-24 10:10:00 28800 3 SSR anomaly
1. Juergen 100427
no bad_data
2010-04-26 11:35:24 19602 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-04-28 18:05:24 18822 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-05-03 04:55:24 24282 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-05-04 03:07:12 29808 1 2 3 4 Sweep interval 15 minutes.
yes TBD
2010-05-05 03:34:30 6 4 Bad spin phase from ISDAT. no bad_data
2010-05-07 13:05:00 40320 1 2 3 4 Sweep interval 15 minutes.
yes TBD
2010-05-09 15:15:24 48282 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-05-12 20:31:34 2 1 Corrupt HX data . no bad_data
2010-05-16 14:40:24 38862 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-05-23 13:02:00 10200 4 SSR controller failure
1. Juergen 1006616
no no_tm
2010-05-27 12:55:24 60282 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-05-28 04:12:14 98580 2 EFW turn-off by DWP watchdog reset.
1. Keith 100529
no no_tm
2010-06-01 00:00:00 -1 3 Bad EDI currents during this month caused occasional extreme (keV) charging on SC3.
no TBD
2010-06-01 13:02:00 36000 3 SSR controller failure. EFW data is patchy and of very suspect quality.
1. Juergen 1006616
no bad_data
2010-06-01 13:05:24 29742 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-06-04 01:00:24 5982 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-06-05 23:40:24 26982 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-06-08 06:36:00 10800 1 HPA /WEC Power Sharing procedure FCP_CLU_M821
1. Colette 100706
yes no_tm
2010-06-08 15:38:00 7200 3 Probe latchup
1. Per-Arne 100608
no bad_data
2010-06-09 23:10:00 22230 2 WEC off by monitoring 444
1. esoc 100611
no bad_data
2010-06-10 07:23:00 10800 1 HPA /WEC Power Sharing procedure FCP_CLU_M821
1. Colette 100706
yes no_tm
2010-06-12 07:42:00 12600 1 HPA /WEC Power Sharing procedure FCP_CLU_M821
1. Colette 100706
yes no_tm
2010-06-14 07:49:00 12540 1 WEC off for HPA in high power mode.
1. Keith 100702
no no_tm
2010-06-16 11:36:00 8400 2 SSR controller failure
1. Juergen 1006617
no no_tm
2010-06-17 06:59:00 8640 1 HPA /WEC Power Sharing procedure FCP_CLU_M821
1. Colette 100714
yes no_tm
2010-06-18 10:35:00 6840 1 HPA /WEC Power Sharing procedure FCP_CLU_M821
1. Colette 100714
yes no_tm
2010-06-21 22:17:00 60 4 Battery crack.
1. Iannis 1006624
no bad_data
2010-06-22 09:55:00 11700 2 Data loss due to STU switchover. AR-171
1. Juergen 1006622
no no_tm
2010-06-22 17:30:24 69882 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-06-22 19:12:00 118800 1 EFW off for unclear reason. Problems with restarting due WECto inrush current. Long data gap.
1. Colette 100715
no no_tm
2010-06-26 16:12:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-06-28 23:50:00 2760 2 Corrupt LX data. HX may or may not contain a burst dump. no bad_data
2010-06-29 04:17:32 122400 3 Long data gap from SSR failure.
1. Juergen 100706
no no_tm
2010-06-30 12:45:24 65082 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-05 18:14:00 4200 1 APPL Buffer Overflow. HX data is bad.
1. Colette 100716
no bad_data
2010-07-05 18:14:24 4122 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-05 18:15:00 3600 2 APPL Buffer Overflow.
1. Colette 100716
no bad_data
2010-07-06 11:53:00 4020 3 WEC monitoring 445.
1. Colette 100716
no bad_data
2010-07-08 00:33:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-10 06:48:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-12 00:40:24 48582 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-14 10:25:24 18882 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-14 16:30:00 300 3 Thruster firing with EFW on (66 pulses) yes bad_data
2010-07-15 06:06:30 75600 2 DWP SEU upset
1. Keith 100802
no no_tm
2010-07-17 01:39:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-17 07:00:00 156600 4 Obscure DWP malfunction.
1. Keith 100729
no no_tm
2010-07-19 07:55:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-20 20:21:03 600 2 Thruster firing with EFW on (124 pulses) yes bad_data
2010-07-21 20:25:24 76282 1 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-21 20:25:24 93582 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-23 19:30:24 7362 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-25 20:30:24 26322 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-28 08:00:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-07-30 01:17:00 53400 2 Bad data. WEC in emergency mode.
1. Keith 100802
no bad_data
2010-07-31 14:45:24 4 3 Bad packet. no bad_data
2010-08-01 15:40:24 25182 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-04 03:50:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-05 15:55:00 1900 2 Perigee raise manoeuvre (440 pulses). yes bad_data
2010-08-06 14:50:00 7200 2 Corrupt data. Cause unknown. Likely commanding error. Recovered with default (-40 nA) bias, so no good data until after next reset at 2010-09-07 18:20. no bad_data
2010-08-06 15:00:00 62460 4 Corrupt data. Cause unknown. Likely commanding error. no bad_data
2010-08-08 00:48:00 6900 1 HPA /WEC Power Sharing procedure FCP_CLU_M821
1. Colette 100903
yes no_tm
2010-08-08 15:34:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-13 03:58:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-14 02:31:00 20400 1 HPA /WEC Power Sharing procedure
1. Colette 100913
yes no_tm
2010-08-15 03:28:00 7200 1 HPA /WEC Power Sharing procedure
1. Colette 100913
yes no_tm
2010-08-15 11:15:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-15 16:32:00 27000 3 WEC off by multiple monitoring entries.
1. Colette 100913
no no_tm
2010-08-16 08:28:21 3600 1 WEC was Power cycle after the CLU_AR-175 (Uncommanded WEC LCL-B On).
1. Colette 100913
no no_tm
2010-08-17 03:30:00 48600 4 No data (LX or HX). Reason unknown. no info
2010-08-20 05:14:36 304200 3 WEC off by multiple monitoring entries. Long data gap while teams asked to analyse WBD temp OOL(3 days).
1. Colette 100913
no no_tm
2010-08-20 23:52:00 7700 1 HPA /WEC Power Sharing procedure
1. Colette 100914
yes no_tm
2010-08-21 10:30:24 69942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-26 12:55:00 387300 4 Problem with EFW-DWP interface. Data lost for 4.5 days until 2010-08-31 00:30.
1. Keith 100827
no bad_data
2010-08-28 14:56:24 40002 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-08-31 16:55:00 27300 4 WEC off by multiple monitoring groups.
1. Jefferys 100901
no no_tm
2010-09-01 15:20:00 24120 3 DWP latchup. Data loss until 2010-09-04 10:20.
1. Keith 100903
no no_tm
2010-09-04 06:39:24 51162 1 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-04 06:39:24 13416 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-09 08:22:24 3942 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-09 22:00:00 66600 1 WEC in emergency mode.
1. Keith 100910
no bad_data
2010-09-11 06:35:24 32082 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-15 08:00:24 73122 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-16 22:18:00 18000 1 HPA /WEC Power Sharing procedure
1. Colette 101001
yes no_tm
2010-09-18 14:30:00 172800 2 WEC problems. 2 days of data lost.
1. Colette 101001
no no_tm
2010-09-20 06:30:24 34782 1 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-20 06:30:24 33336 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-20 06:30:24 34782 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-25 17:35:43 2 1 Bad packet. no bad_data
2010-09-26 01:02:23 3 3 Bad packet. no bad_data
2010-09-27 05:18:24 23142 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-09-28 21:44:00 14400 1 HPA /WEC Power Sharing procedure
1. Colette 101014
yes no_tm
2010-10-02 00:20:00 65400 2 WEC off by onboard monitor.
1. Colette 101014
no no_tm
2010-10-02 21:16:00 18000 1 HPA /WEC Power Sharing procedure
1. Colette 101014
yes no_tm
2010-10-04 04:28:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-10-10 06:24:24 68502 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-10-10 21:04:40 5 4 Bad phase data. no bad_data
2010-10-16 22:06:59 2 3 Bad packet. no bad_data
2010-10-17 06:10:24 50682 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-10-17 15:26:50 3 3 Bad packet. no bad_data
2010-10-19 23:20:00 45000 2 SSR latchup. CLU_AR-177.
1. Juergen 101026
no no_tm
2010-10-20 08:09:00 12600 2 EFW switched off for unknown reason.
1. Per-Arne 101021
no no_tm
2010-10-23 16:30:00 80100 1 EFW probe latchup and telemetry problems.
1. Per-Arne 101025
no bad_data
2010-10-24 19:30:00 162900 1 Bad biasing on P4, looks like other problems, too. Very bad data.
1. Per-Arne 101025
2. Per-Arne 101026
no bad_data
2010-10-26 18:14:00 10800 3 CTU switchover. CLU_AR-179
1. Silvia 101028
no no_tm
2010-10-27 00:57:00 23280 4 CTU switchover. CLU_AR-180
1. Silvia 101028
no no_tm
2010-10-29 01:36:24 7422 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-10-29 08:14:24 97722 1 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-10-30 04:27:24 24942 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-02 23:40:00 88800 3 Probe latchups, then a DWP upset
1. Keith 101104
no bad_data
2010-11-07 04:15:24 4002 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-07 11:10:00 73800 1 Choppy telemetry, digital errors.
no bad_data
2010-11-08 11:10:00 43200 2 WEC off by monitoring 444/445/446.
1. Keith 101109
no no_tm
2010-11-08 13:45:24 78702 1 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-08 13:45:24 34356 2 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-08 13:45:24 78702 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-11 02:16:24 56202 1 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-18 02:05:24 30882 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-11-18 21:00:00 28140 4 Data loss due to ground segment problem.
1. AR-2733
no no_tm
2010-11-22 02:10:00 64800 3 WEC off by monitoring 444.
1. Colette 101224
no no_tm
2010-11-28 12:09:00 25200 3 WEC off by multiple monitoring entries.
1. Colette 101230
no no_tm
2010-12-10 22:40:24 17442 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2010-12-13 05:05:24 16242 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-02-26 14:16:24 62802 1 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-03-05 18:13:40 -1 3 WEC switched off by spacecraft monitoring and could not be switched on again by standard procedures. Investigations identified likely cause and led to tests in late May. If successful, operations could possibly be resumed in June.
1. Meeting ESOC 110513
no no_tm
2011-03-11 06:45:24 14082 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-03-31 05:04:24 36342 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-04-08 20:38:00 14400 2 To be investigated by AIE no bad_data
2011-04-21 09:00:00 769414 2 Probe contamination at extremely low perigee passes caused P2 data to become intolerably bad from about this time (perigee at 09:35) until the problem was discovered, analyzed, tests were run in space, and the bias current finally was adjusted from 2011-04-30 06:43:34.
no bad_bias
2011-04-24 03:25:24 65442 1 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-04-28 11:07:24 79542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-04-30 06:43:31 -1 2 The adjusted bias current for low photoemission current due to probe contamination at low perigee passes should be effective from this time, and data should thus be of acceptable quality.
1. Per-Arne 110429
yes info
2011-05-02 02:45:24 24882 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-05-06 20:16:24 21222 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-05-22 17:15:00 19200 3 4 Sweeps also on SC34, 30 min interval.
yes TBD
2011-05-26 12:05:00 660 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-05-31 00:35:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-01 09:31:24 -1 3 EFW operational again after WEC on SC3 was brought back to life. However, P3 does not work any longer after this date.
yes no_p3
2011-06-04 12:29:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-09 01:38:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-13 14:55:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-16 19:18:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-22 16:08:00 720 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-25 18:19:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-06-27 03:45:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-03 22:31:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-04 21:19:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-06 07:30:24 74862 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-07-11 15:19:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-13 18:34:24 7542 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-07-15 02:19:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-18 08:19:24 8442 1 2 3 18 kHz internal burst rate. SPLIT sampling table caused bad p32 data
1. SPLIT mode problem
no no_p32
2011-07-18 10:19:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-22 00:42:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-24 06:58:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-07-31 01:35:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-08-06 20:14:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-08-13 14:49:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-08-20 09:27:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-08-22 13:11:51 279240 2 Bad or non-existing EFW data. Exact reason not known, but the problem started following a DWP anonaly.
1. Per-Arne 110822
2. Keith 110831
no bad_tm
2011-08-27 04:13:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-09-05 05:23:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-09-12 00:12:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval.
yes TBD
2011-10-09 03:54:30 -1 1 3 4 Bias current changed to 140 nA on SC134 as solar UV flux has increased.
yes info
2011-10-29 12:15:36 60 1 4 Bias stepping from 140 to 320 nA in steps of 12 nA, 4 sec step rate.
yes bad_bias
2011-10-30 02:49:00 52380 2 WEC turned off by onboard monitoring.
1. Keith 111031
no TBD
2011-10-30 10:01:00 900 1 3 4 Sweeps on all s/c, 64 s interval, with 450 Hz burst collection by time-tagged trig just after start. Burst quantities (same for all s/c): V1V2M, V3V4M, V1V2U, V3V4U. (Was planned also for C2, but lost due to WEC turned off on that s/c by the problem noted above).
yes TBD
2011-10-31 21:30:24 3600 1 3 4 Special burst targeting narrowband noise.
1. Anders 111011
yes TBD
2011-10-31 22:05:24 1500 2 Special burst targeting narrowband noise.
1. Anders 111011
yes TBD
2011-11-03 03:45:24 3600 1 2 4 Special burst targeting narrowband noise.
1. Anders 111011
yes TBD
2011-11-03 04:31:24 480 3 Special burst targeting narrowband noise.
1. Anders 111011
yes TBD
2011-11-04 03:42:00 97980 3 Bad WEC telemetry, DWP buffer overflow? no bad_tm
2011-12-04 00:10:00 136 4 Bias stepping P34 from 8 to 380 nA in steps of 12 nA, 4 sec step rate. Start by setting 8 nA, 4 nA, 8 nA (to give time for any transient).
yes bad_bias
2011-12-04 00:10:00 136 2 Bias stepping P2 from 4 to 132 nA in steps of 4 nA, 4 sec step rate.
yes bad_bias
2011-12-05 15:15:00 900 1 2 3 4 Sweeps on all s/c, 64 s interval, with 450 Hz burst collection by time-tagged trig just after start. Burst quantities (same for all s/c): V1V2M, V3V4M, V1V2U, V3V4U.
yes TBD
2012-01-23 10:51:19 78600 1 WEC data lost following DWP watchdog reset.
1. Keith 120124
no no_tm
2012-02-03 22:08:00 300 1 2 3 4 Sweeps during BM3 interval that could not be used for dumping internal burst. Sweeps start at 22:01 and go on into the BM3.
yes TBD
2012-02-11 19:05:00 85700 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-14 00:35:00 88200 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-16 09:05:00 99000 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-18 14:35:00 80800 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-20 19:35:00 87000 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-23 03:05:00 83100 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-25 09:05:00 77100 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-27 15:05:00 85200 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-02-29 21:35:00 80400 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-03-03 05:35:00 75500 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-03-04 02:51:00 280 2 Bias stepping (slow sweep) on P2 (first vbias steps of 10 V from -39 V to -19.4 V, then vbias steps of +0.6 V from -19.4 V to -0.6 V, then ibias steps of -4 nA from -12 nA to -140 nA).
1. Anders 120214
yes bad_bias
2012-03-04 02:54:00 280 4 Bias stepping (slow sweep) on P34 (first vbias steps of 10 V from -39 V to -19.4 V, then vbias steps of +0.6 V from -19.4 V to -0.6 V, then ibias steps of -12 nA from -104 nA to -500 nA).
1. Anders 120214
yes bad_bias
2012-03-04 02:55:40 140 2 Bias stepping (slow sweep) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Stepping on P2 starts at 02:51.
1. Anders 120214
yes bad_bias
2012-03-05 10:05:00 84600 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-03-07 17:35:00 80400 2 P2 at maximum positive bias voltage +39 V.
yes no_p2
2012-03-14 03:09:00 140 2 Bias voltage stepping (slow sweep) on P2. Modified from 120304.
1. Anders 120214
yes no_p2
2012-03-14 03:09:00 140 4 Bias voltage stepping (slow sweep) on P34. Modified from 120304.
1. Anders 120214
yes no_p34
2012-03-14 03:11:20 140 2 Bias current stepping (slow sweeps) on P2 (ibias steps of -4 nA from -12 nA to -140 nA). Modified from 120304.
1. Anders 120214
yes bad_bias
2012-03-14 03:11:20 140 4 Bias current stepping (slow sweeps) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Modified from 120304.
1. Anders 120214
yes bad_bias
2012-03-14 03:13:40 140 2 Bias stepping (slow sweeps) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Modified from 120304.
1. Anders 120214
yes bad_bias
2012-04-10 04:29:00 140 2 Bias voltage stepping (slow sweep) on P2. Operation identical to 120314.
1. Anders 120214
yes no_p2
2012-04-10 04:29:00 140 4 Bias voltage stepping (slow sweep) on P34. Operation identical to 120314.
1. Anders 120214
yes no_p34
2012-04-10 04:31:20 140 2 Bias current stepping (slow sweeps) on P2 (ibias steps of -4 nA from -12 nA to -140 nA). Operation identical to 120314.
1. Anders 120214
yes bad_bias
2012-04-10 04:31:20 140 4 Bias current stepping (slow sweeps) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Operation identical to 120314.
1. Anders 120214
yes bad_bias
2012-04-10 04:33:40 140 2 Bias stepping (slow sweeps) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Operation identical to 120314.
1. Anders 120214
yes bad_bias
2012-06-10 06:49:00 140 2 Bias voltage stepping (slow sweep) on P2. Operation identical to 120314 and 120410.
1. Anders 120214
yes no_p2
2012-06-10 06:49:00 140 4 Bias voltage stepping (slow sweep) on P34. Operation identical to 120314 and 120410.
1. Anders 120214
yes no_p34
2012-06-10 06:51:20 140 2 Bias current stepping (slow sweeps) on P2 (ibias steps of -4 nA from -12 nA to -140 nA). Operation identical to 120314 and 120410.
1. Anders 120214
yes bad_bias
2012-06-10 06:51:20 140 4 Bias current stepping (slow sweeps) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Operation identical to 120314 and 120410.
1. Anders 120214
yes bad_bias
2012-06-10 06:53:40 140 2 Bias stepping (slow sweeps) on P34 (ibias steps of -4 nA from -12 nA to -140 nA). Operation identical to 120314 and 120410.
1. Anders 120214
yes bad_bias
2012-08-16 20:09:40 347141 1 All probes latched at -68 V, reason unknown.
1. Keith 120820
no bad_data
2012-08-20 20:35:21 46939 1 Probe 2 latched at +68 V, reason unknown. no no_p2
2012-09-04 06:40:00 140 1 Bias current stepping (slow sweep) on P2. yes bad_bias
2012-09-04 06:40:00 140 2 Bias voltage stepping (slow sweep) on P2. yes no_p2
2012-09-04 06:40:00 140 3 Bias current stepping (slow sweep) on P2. yes bad_bias
2012-09-04 06:40:00 140 4 Bias current stepping (slow sweep) on P12. yes bad_bias
2012-09-04 06:42:20 140 1 Bias current stepping (slow sweep) on P3. yes bad_bias
2012-09-04 06:42:20 140 2 Bias current stepping (slow sweep) on P2. yes bad_bias
2012-09-04 06:42:20 140 3 Bias current stepping (slow sweep) on P4. yes bad_bias
2012-09-04 06:42:20 140 4 Bias current stepping (slow sweeps) on P34. yes bad_bias
2012-09-04 06:44:40 140 2 Bias current stepping (slow sweeps) on P34. yes bad_bias
2012-09-04 06:44:40 1946 4 P3 bias current test. yes bad_bias
2012-09-04 06:47:00 1786 2 P2 and P4 bias current test. yes bad_bias
2012-09-04 06:47:00 1786 2 P2 and P4 bias current test. yes bad_bias
2012-09-10 16:48:00 720 1 4 WBD interference test (guard stepping).
1. Anders 120817
yes bad_bias
2012-10-14 02:13:30 604800 1 2 3 4 Guard voltage changed to -2 V on all working probes for a week.
yes info
2012-10-20 21:11:30 604800 1 2 3 4 Guard voltage changed to -8 V on all working probes for a week.
yes info
2012-10-27 16:12:30 -1 1 2 3 4 Guard voltage changed to -4 V on all working probes permanently.
yes info
2013-07-01 14:43:44 -1 4 Probe 4 effectively lost from this time. no no_p4
2014-10-03 06:20:20 32916 2 EFW reset with launch default values.
1. Per-Arne 150923
no bad_data
2014-10-07 21:03:00 268620 2 DWP problem causing bad or no EFW telemetry.
1. Per-Arne 150923
no bad_data
2014-10-15 23:54:00 54360 2 Probes saturated, possibly due to low density.
1. Per-Arne 150923
no bad_data
2014-10-23 15:11:00 159240 3 Bad EFW data after incorrect startup.
1. Per-Arne 150923
no bad_data
2014-10-25 00:53:00 41832 2 Probes saturated, possibly due to low density.
1. Per-Arne 150923
no bad_data
2014-10-31 09:46:08 49944 4 DWP problem causing bad EFW data.
1. Per-Arne 150923
no bad_data
2014-11-01 15:05:20 78788 2 Probes saturated, possibly due to low density, followed by DWP problem.
1. Per-Arne 150923
no bad_data
2014-11-03 20:58:17 -1 3 Probe 2 effectively lost from this time. no no_p2
2014-11-15 08:36:00 64500 1 Bad EFW data after spontaneous EFW reset.
1. Per-Arne 150923
no bad_data
2015-02-15 02:12:00 17880 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-02-17 07:36:30 -1 4 Probe 3 effectively lost from this time. no no_p3
2015-02-23 11:30:00 6600 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-02-27 20:54:00 12000 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-02-28 14:10:00 10200 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-09 09:19:26 50 1 Maneuvre. yes bad_data
2015-03-15 16:54:00 720 1 2 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-15 17:42:00 9180 1 2 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-16 10:40:00 14880 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-17 14:07:33 1048 1 Maneuvre. yes bad_data
2015-03-20 17:00:00 88800 1 Probe 3 latched at -10.29 V. no no_p3
2015-03-31 13:47:45 540 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-31 14:11:30 95 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-31 14:54:30 120 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-03-31 17:14:00 540 1 2 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-02 20:03:42 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-02 20:14:39 40 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-02 20:29:30 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-02 21:27:10 76 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-09 12:37:00 3180 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 06:40:22 57 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 06:52:00 6660 1 2 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 09:10:00 2160 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 11:25:00 300 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 12:34:00 240 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 12:36:22 32 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-10 16:58:00 480 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-11 16:39:20 80 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-11 16:46:25 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-11 21:45:00 300 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-12 16:47:00 7560 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-12 19:42:12 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-12 21:52:48 16 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-13 23:56:32 59 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:02:40 70 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:05:58 262 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:14:30 400 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:23:14 53 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:29:38 118 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:30:00 223 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:38:00 330 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 00:52:00 7680 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:10:58 137 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:11:50 95 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:17:52 413 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:20:30 125 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:28:45 70 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:29:20 39 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:31:21 69 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:34:20 17950 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 19:34:42 9193 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 22:12:30 248 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 22:24:12 70 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 22:26:18 20 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 23:17:14 191 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 23:31:22 1404 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-14 23:56:06 482 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-15 00:29:38 52 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-15 00:38:30 1210 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-15 00:39:58 67 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-15 00:52:34 228 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-15 01:07:15 205 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-16 20:16:10 284 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-16 20:22:13 86 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-20 20:09:20 25 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-20 20:24:22 10 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-20 20:25:32 26 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-20 21:08:00 191 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-20 21:13:30 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-20 21:18:49 144 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 09:18:52 23 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 09:23:45 13 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 09:26:30 189 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 10:10:34 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 10:12:54 9 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 10:52:55 45 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-26 10:59:05 13 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-27 07:59:48 11 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-27 08:30:01 29 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-27 11:25:26 124 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-27 11:54:42 657 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-27 14:05:37 3514 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-27 15:04:38 10 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-28 18:03:55 865 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-28 19:01:38 468 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 12:58:45 58 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 13:00:17 14 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 13:28:25 10 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 13:35:50 53 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 13:44:29 164 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 13:56:42 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:05:12 43 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:09:23 71 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:12:25 249 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:22:05 354 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:30:10 25 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:33:30 210 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:43:55 30 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:49:14 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:50:42 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:51:46 168 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 16:57:50 2316 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 17:21:58 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 17:43:50 1345 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 21:05:15 540 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 21:19:14 3331 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-04-29 22:23:33 13 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 10:42:46 9 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 10:44:01 49 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 12:44:02 49 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 14:52:10 110 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 15:02:40 950 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 15:16:10 140 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 15:27:20 3115 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:00:18 152 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:05:15 160 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:05:30 149 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:09:50 137 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:09:55 133 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:21:25 14 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:21:25 281 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:23:30 161 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:30:18 232 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:30:26 216 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:36:26 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:41:26 144 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 16:49:38 629 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 17:05:22 23 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 17:57:54 88 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 18:00:51 91 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-06 18:07:58 542 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 17:13:10 4723 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 17:55:30 45 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:08:25 130 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:25:27 45 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:28:25 23 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:35:42 144 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:46:41 941 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:48:10 35 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:51:52 78 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 18:57:26 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 19:04:45 657 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 19:04:57 101 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 20:15:30 24 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 20:31:06 1474 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-09 20:34:14 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-10 12:17:17 37 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-10 15:08:50 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-10 16:31:53 110 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-11 23:42:55 128 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-11 23:49:12 38 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-11 23:51:17 2863 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 00:42:12 38 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 00:47:45 295 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 01:03:17 5601 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 01:26:14 41 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 02:03:27 40 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 02:08:17 106 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 02:12:44 179 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 02:23:58 76 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 02:26:45 365 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-12 04:24:28 4466 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 15:53:30 804 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:02:57 45 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:04:38 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:14:38 6280 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:15:22 46 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:24:10 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:25:54 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:27:37 21 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:29:20 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:38:18 16 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:41:42 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:44:10 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:50:42 10 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:51:30 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 16:54:26 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:00:46 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:02:14 16 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:03:30 14 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:06:54 26 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:07:50 16 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:08:30 28 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:09:38 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:16:14 20 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:19:42 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:20:06 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:20:42 16 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:22:46 10 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:44:10 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 17:56:02 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 18:10:10 533 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 18:24:38 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 18:38:20 18 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 18:40:05 25 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 18:41:50 41 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 19:41:34 656 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 19:41:58 20 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 19:51:11 24 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 20:35:21 13 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 20:49:01 684 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 20:50:46 9 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 21:01:34 68 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 21:24:51 11 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 23:35:50 256 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-23 23:58:22 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 00:38:25 21 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 02:27:10 15 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 03:08:10 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 03:37:32 798 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 04:00:08 11 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 04:02:02 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 08:32:34 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 08:37:57 21 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 10:54:24 66 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 11:00:45 80 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 11:03:08 11 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 11:33:40 42 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 11:38:30 10 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 11:42:50 45 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-24 12:04:24 42 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-25 21:18:40 1400 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-25 21:37:12 1474 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-25 22:16:40 6166 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 00:00:50 11221 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 03:16:13 1186 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 04:10:30 145 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 06:24:58 40832 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 15:07:46 1888 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 15:48:05 1759 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 16:33:50 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 16:59:30 820 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 17:34:40 505 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-26 17:50:40 134 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-27 21:48:50 690 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-27 22:31:21 153 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-27 23:15:13 2677 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 00:01:22 3284 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 01:37:14 1656 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 02:12:42 1133 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 02:37:05 82 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 02:47:35 715 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 03:14:46 840 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 04:05:54 23652 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 10:45:30 324 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 11:13:10 532 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 11:30:18 96 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 11:41:30 170 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 11:46:18 156 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 11:52:58 532 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 12:07:38 1200 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 12:31:26 1579 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 12:57:22 12 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 13:14:34 38536 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 13:48:05 19 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 13:59:22 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 14:02:38 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 14:09:10 17 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-28 14:11:45 19 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 00:25:22 43 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 01:04:08 1552 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 01:31:18 327 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 01:56:45 1425 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 02:27:40 56 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 02:36:41 233 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-29 17:09:11 95 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-30 05:41:34 13 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-05-30 15:41:34 16 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 00:27:29 41 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 01:05:45 165 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 01:48:20 380 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 02:01:14 166 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 03:10:10 29 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 04:36:06 36 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 04:55:15 127 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-04 05:26:56 25 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 00:06:35 16579 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 03:41:14 2064 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 04:35:45 389 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 04:53:35 285 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 04:53:54 276 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 04:59:22 448 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 05:01:41 181 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 05:06:06 8 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 11:32:50 21416 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 18:06:00 543 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 19:10:42 8020 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 22:05:17 133 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 22:38:46 41 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 22:41:30 47 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 23:02:50 80 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 23:07:58 96 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-06 23:17:54 68 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-08 01:09:50 356 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-08 01:21:45 57 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-08 05:50:22 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-12 16:03:45 12822 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-12 17:02:05 234 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-12 17:32:22 123 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-12 17:38:10 28 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-12 19:43:14 261 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-21 16:48:32 28848 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-21 16:50:14 11369 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-21 20:24:16 6951 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-22 05:59:46 28192 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-22 14:42:34 346 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-22 14:55:16 590 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-25 13:29:15 7 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-25 13:31:14 16 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-25 13:31:50 52 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-25 13:33:54 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-25 13:35:02 28 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-30 17:46:54 17260 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-30 20:43:39 69 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-30 23:34:38 140 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-30 23:44:54 46 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-30 23:47:22 41 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-06-30 23:59:42 28 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 00:04:46 564 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 00:44:57 25 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 02:10:44 894 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 02:36:22 216 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 02:43:25 130 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 02:48:50 204 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 03:02:05 1625 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 03:31:42 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 03:34:10 80 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-01 03:37:25 210 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 00:11:50 6780 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 00:58:30 3603 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 02:07:33 41 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 02:20:26 194 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 02:34:18 6980 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 04:34:51 243 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 04:42:34 1044 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 05:03:55 413 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 05:17:58 17612 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 10:17:26 1168 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 10:42:20 946 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 10:55:34 3728 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 12:07:38 256 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 12:25:30 184 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 12:37:34 4320 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 13:45:02 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 13:48:42 416 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 13:59:10 292 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 14:09:18 192 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-07-21 14:34:02 168 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-07 14:15:10 642 1 2 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-07 20:22:30 886 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-07 20:23:40 394 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-12 07:24:25 512 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-12 07:33:30 260 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-14 15:22:15 51 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-14 15:51:09 217 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-15 08:32:58 111 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-15 08:53:45 126 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-15 09:22:30 1685 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-15 09:54:35 2099 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-15 10:34:16 835 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-15 11:22:26 1009 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-16 07:41:46 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-16 08:07:02 528 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-16 18:23:10 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-16 18:23:50 56 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-16 19:44:08 65 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-16 20:57:30 109 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-18 22:04:06 166 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 01:42:42 38 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 01:55:46 206 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 01:59:02 138 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 02:05:42 336 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 02:44:50 393 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 03:07:50 345 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 07:05:38 884 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 07:28:18 248 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 07:39:42 280 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 08:10:26 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 08:41:38 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 08:43:02 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 08:43:37 13 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-19 08:45:06 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-20 22:37:54 38 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-08-30 12:06:48 2862 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-06 10:38:28 7 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-06 11:04:29 6 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-08 01:27:30 557 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-08 11:38:56 1263 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-11 01:15:10 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-21 06:52:09 20596 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-09-26 14:14:14 13281 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-05 12:53:03 34 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 10:05:42 10 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 12:23:50 412 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 12:51:10 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 13:11:46 256 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 13:54:14 808 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 14:19:50 349 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 15:38:26 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 15:41:38 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 15:42:14 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 16:53:30 32 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 16:54:30 32 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-07 17:00:02 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-10-30 16:12:22 536 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-02 23:30:00 558 1 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-03 00:51:54 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-03 01:02:14 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-03 01:04:30 25188 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-03 01:38:30 23146 1 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-07 00:40:46 984 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-10 10:49:02 344 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-10 11:02:01 368 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-17 17:53:14 68 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-19 10:52:22 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-27 10:27:06 1224 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-27 11:18:18 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-27 11:21:54 1948 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-27 12:11:42 116 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-27 12:59:38 60913 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-28 08:57:58 16192 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-28 11:35:29 3237 1 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-28 20:27:14 61 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-28 20:36:16 4742 1 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-29 18:42:50 54088 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-29 19:42:45 105 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-29 19:51:00 651 1 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-29 20:24:06 419 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-29 21:28:41 199 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 00:01:42 1218 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 00:32:30 260 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 01:45:10 644 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 02:11:42 1209 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 02:57:26 545 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 03:49:02 14492 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 08:10:14 185 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 08:33:49 269 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-11-30 08:44:41 85 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-05 08:47:22 989 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-05 09:13:46 48 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-05 19:24:58 656 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-07 21:49:01 66 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-08 00:48:25 437 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-10 08:09:14 1932 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 15:42:42 1668 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 16:43:32 217 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 16:53:29 349 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 17:44:06 849 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 17:44:06 849 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 18:13:10 486 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 18:29:54 635 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 18:33:18 32 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 18:47:42 64 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 18:50:54 699 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-14 19:07:02 3532 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-19 18:03:42 21903 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-19 18:05:38 21192 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 00:26:58 3664 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 00:35:34 1700 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 03:17:30 10540 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 06:26:30 49 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 13:55:46 1372 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 14:43:42 104 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 14:51:14 10972 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 15:06:34 91 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-20 21:54:18 312 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-22 07:00:15 769 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-22 10:16:26 2764 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 13:12:34 10 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 17:13:22 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 17:41:30 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 18:16:50 8 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 18:32:34 83 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 19:56:50 10 1 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 19:58:04 74 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-27 20:02:06 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-28 08:02:34 865 2 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-28 10:35:37 193 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 02:00:18 27 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 02:26:54 46 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 08:40:42 164 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 08:48:14 16 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 08:53:14 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 08:53:14 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 09:08:22 366 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 10:23:18 728 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 11:00:47 3371 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 12:51:22 1912 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 14:51:34 20 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 14:55:42 12 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 15:17:58 28 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 17:09:50 2499 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-29 18:08:38 96 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-30 02:49:29 245 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-30 03:01:42 567 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-30 03:47:46 648 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-30 04:13:50 48 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-30 04:25:02 268 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-31 05:34:05 1722 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-31 06:26:18 4692 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-31 08:33:37 597 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-31 08:33:37 597 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-31 09:55:01 3301 4 Probes saturated in dense plasma due to high bias. no bad_data
2015-12-31 13:44:10 2384 4 Probes saturated in dense plasma due to high bias. no bad_data